### v1

# Moving sum procedure for multiple change point detection in large factor models

Abstract
The paper proposes a moving sum methodology for detecting multiple change points in high-dimensional time series under a factor model, where changes are attributed to those in loadings as well as emergence or disappearance of factors. We establish the asymptotic null distribution of the proposed test for family-wise error control, and show the consistency of the procedure for multiple change point estimation. Simulation studies and an application to a large dataset of volatilities demonstrate the competitive performance of the proposed method.
11footnotetext: Department of Economics, Università di Bologna. Email: matteo.barigozzi@unibo.it.22footnotetext: School of Mathematics, University of Bristol. Email: haeran.cho@bristol.ac.uk.33footnotetext: Department of Economics and Management, Universita’ di Pavia and School of Business, University of Leicester. Email: lt285@leicester.ac.uk.
Keywords: data segmentation, MOSUM, factor model
MSC subject classification: 62M10, 62H25
1 Introduction
Factor models are, arguably, one of the most frequently employed tools to model and carry out inference when large-dimensional, vector-valued time series are available. Whilst a comprehensive review of the literature goes well above and beyond the scope of this paper, it is worth noting that factor models have a long history, dating back at least to the seminal paper by Spearman, (1904). Since being popularised by the contribution by Chamberlain and Rothschild, (1983), and since the development of the asymptotic theory to analyse large dimensional factor models (Bai,, 2003) their usage has become de rigueur in social sciences and economics where they have been applied to diverse fields such as business cycle analysis, asset pricing and economic monitoring and forecasting (see, inter alia, the review by Stock and Watson, (2011) for a comprehensive list of references).
As the time dimension increases, it is inevitable that any model (including factor models) may undergo changes in its structure which, in turn, may affect the properties of estimation techniques and hamper predictive ability. There is a huge literature on both ex-post and online detection of change points in various contexts, and we refer to the recent book by Horváth and Rice, (2024) for a review focussed in particular on ex-post detection, and to the paper by Aue and Kirch, (2024) for a review of the literature on online (sequential) detection.
When factor models are concerned, on the one hand, the analysis in Stock and Watson, (2009) and Bates et al., (2013) shows that, when such changes are “sufficiently small” and do not involve a change in the number of common factors, the presence of change points is inconsequential for inference on the factor spaces. On the other hand, in several applications, there is no guarantee that the factor structure only undergoes small changes and that the number of common factors is time-invariant. Hence, it is not surprising that the literature has developed several methods to test, retrospectively, whether there is a change point or not. Breitung and Eickmeier, (2011), Corradi and Swanson, (2014), Han and Inoue, (2015), Yamamoto and Tanaka, (2015), Baltagi et al., (2017), Cheng et al., (2016), Duan et al., (2023), and, recently, Bai et al., (2024) propose several tests to check if there is a break, based on the idea that a change point in the loadings is equivalent to a change in the covariance matrix of the common factors. Thus assuming homoscedastic factors, a change in the covariance matrix of the estimated common factors can only be due to a break in the loadings. See also Massacci, (2023) and, in an online context, Barigozzi and Trapani, (2020) and He et al., (2024) for alternative formulation of change point tests.
Whilst, on account of the references above, the literature on testing for change points seems well-established, there is relatively little work when the estimation of a change point is concerned. In particular, the literature largely focuses on the At-Most-One-Change alternative (Duan et al.,, 2023), and few addresses the detection and estimation of (possibly) multiple change points in the factor structure, with some exceptions (Barigozzi et al.,, 2018; Li et al.,, 2023).
In this paper, we make an advance on the change point literature by proposing a method to estimate the number and locations of (possibly) multiple change points under a (static) factor model. Specifically, we propose a procedure based on the so-called MOving SUMs (MOSUM) process which, since the seminal contribution by Hušková and Slabỳ, (2001), has been shown to have desirable properties in multiple change point estimation under weak assumptions on the distribution of the data. Versatility of the MOSUM procedure has been demonstrated in the context of detecting changes in the mean (Hušková and Slabỳ,, 2001; Eichinger and Kirch,, 2018), trend (Kim et al.,, 2024) and the drift of stochastic processes (Kirch and Klein,, 2024); see also Kirch and Reckruehm, (2024) for a general framework for multivariate time series segmentation based on an estimating function. There also exist recent contributions which extend the use of this methodology to high-dimensional time series (Cho and Owens,, 2024; Cho et al.,, 2024), demonstrating its scalability and suggesting that it is worth exploring its performance in the context of large factor models.
We build on that, as discussed above, a change in the loadings is observationally equivalent to a change in the covariance matrix of the common factors, and construct a MOSUM statistic based on the (moving) partial sums of the outer products of the estimated factors. In this respect, our statistic is akin to the one based on the (maximally selected) sequence of likelihood ratio tests proposed e.g. in Duan et al., (2023), save for the fact that it is based on using moving, as opposed to cumulative, partial sums. Upon testing for changes in the factor model using the MOSUM process, if the null hypothesis of no changes is rejected, we perform multiple change point detection by estimating both the total number and the locations of the breaks. We establish the asymptotic null distribution of the MOSUM process as well as consistency in multiple change point detection, with the accompanying rate of estimation for individual breaks. In doing so, we derive a Strong Invariance Principle for the MOSUM process based on the outer products of the estimated factors, which may be of independent interest.
The remainder of the paper is organised as follows. We describe the MOSUM procedure in Section 2. Assumptions, the limiting distribution under the null, and the consistency under the alternative are reported in Section 3. Sections 4 and 5 provide findings from simulation studies and an application to financial data. Section 6 concludes the paper, and all the proofs of theoretical results are given in Appendix. An implementation of the proposed method is available at https://github.com/haeran-cho/mosum.fts.
Notations.
For a random variable $X$, we have $|X|_{\nu}=(\mathsf{E}(|X|^{\nu}))^{1/\nu}$. We define: weak convergence on the space $\mathcal{D}[0,1]$ with $\overset{\mathcal{D}}{\rightarrow}$; =$\overset{\mathcal{D}}{=}$ is equality in distribution; $\rightarrow$ is the ordinary limit; $\Gamma(\cdot)$ is the Gamma function; $\log(x)$ is the natural logarithm of $x>0$; $\mathbb{I}_{\mathcal{A}}$ is an indicator function satisfying $\mathbb{I}_{\mathcal{A}}=1$ if the event $\mathcal{A}$ is true and $\mathbb{I}_{\mathcal{A}}=0$ otherwise. By $\mathbf{I}$, $\mathbf{O}$ and $\mathbf{0}$, we denote an identity matrix, a matrix of zeros and a vector of zeros whose dimensions depend on the context. For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, we denote by $\mathbf{A}^{\top}$ its transpose and $\|\mathbf{A}\|$ its Euclidean norm, with $\Lambda_{\max}(\mathbf{A})$ and $\Lambda_{\min}(\mathbf{A})$ denoting its largest and the smallest eigenvalues in modulus. When $m=n$, we denote by $\mathsf{Vech}(\mathbf{A})$ the vector of length $m(m+1)/2$ that stacks the elements on and below the main diagonal of $\mathbf{A}$, and $\mathbf{L}_{m}$ the $m(m+1)/2\times m^{2}$-matrix satisfying $\mathsf{Vech}(\mathbf{A})=\mathbf{L}_{m}\mathsf{Vec}(\mathbf{A})$, where $\mathsf{Vec}(\cdot)$ is an operator that stacks the columns of the matrix into a vector. Conversely, we define $\mathbf{K}_{m}\in\mathbb{R}^{m^{2}\times m(m+1)/2}$ satisfying $\mathsf{Vec}(\mathbf{A})=\mathbf{K}_{m}\mathsf{Vech}(\mathbf{A})$, see Appendix A.12 of Lütkepohl, (2005). Finally, given two sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ if, for some finite positive constant $C$ there exists $N\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ such that $|a_{n}||b_{n}|^{-1}\leq C$ for all $n\geq N$.
2 MOSUM procedure for data segmentation
2.1 Model
[…]
$\mathbf{X}_{t}=\left\{\begin{array}[]{ll}\bm{\Lambda}_{0}\mathbf{f}_{t}+\mathbf{e}_{t}&\text{for \ }k_{0}+1=1\leq t\leq k_{1},\\ \bm{\Lambda}_{1}\mathbf{f}_{t}+\mathbf{e}_{t}&\text{for \ }k_{1}+1\leq t\leq k_{2},\\ \vdots&\\ \bm{\Lambda}_{R}\mathbf{f}_{t}+\mathbf{e}_{t}&\text{for \ }k_{R}+1\leq t\leq k_{R+1}=T.\end{array}\right.$ (1)
Here, the matrix of loadings $\bm{\Lambda}_{j}=[\bm{\lambda}_{j,1},\ldots,\bm{\lambda}_{j,N}]^{\top}\in\mathbb{R}^{N\times r}$ has column rank $r_{j}\leq r$ fixed for all $N$, and “loads” the vector of random factors $\mathbf{f}_{t}=(f_{1,t},\ldots,f_{r,t})^{\top}$ onto the cross-sections of $\mathbf{X}_{t}$, and $\mathbf{e}_{t}=(e_{1,t},\ldots,e_{N,t})^{\top}$ denotes the idiosyncratic component. The model in (1) allows for changes due to emergence or disappearance of factor(s) by permitting the ranks of $\bm{\Lambda}_{j}$ to vary, as well as rotational changes in the loading matrices. At the same time, the model is not identifiable in that the changes in the loadings are (observationally) equivalent to changes in the second-order properties of the common factor $\mathbf{f}_{t}$ – a fact, as mentioned in Introduction, frequently explored in the relevant literature for developing change point tests in the At-Most-One-Change scenario; see e.g. Bai et al., (2024). We further illustrate this point in the following example.
Example 1.
Consider the factor model in (1) with $R=1$ and the single change point at $k_{1}=k^{\ast}$. For any $\bm{\Lambda}_{j}\in\mathbb{R}^{N\times r_{j}}$, $j=0,1$, we can find $\bm{\Lambda}\in\mathbb{R}^{N\times r}$ of full column rank with $r\geq\max(r_{0},r_{1})$ such that $\bm{\Lambda}_{j}=\bm{\Lambda}\mathbf{A}_{j}$ for $\mathbf{A}_{j}\in\mathbb{R}^{r\times r}$ of rank $r_{j}$. Then, we can re-write (1) as
$\begin{bmatrix}\mathbf{X}_{0:k^{\ast}}\\ \mathbf{X}_{k^{\ast}:T}\end{bmatrix}=\begin{bmatrix}\mathbf{F}_{0:k^{\ast}}\bm{\Lambda}_{0}^{\top}\\ \mathbf{F}_{k^{\ast}:T}\bm{\Lambda}_{1}^{\top}\end{bmatrix}+\mathbf{E},\text{ \ where \ }\mathbf{E}^{\top}=\begin{bmatrix}\mathbf{e}_{1},\ldots,\mathbf{e}_{T}\end{bmatrix},$
$\underset{(b-a)\times N}{\mathbf{X}_{a:b}}=\begin{bmatrix}\mathbf{X}_{a+1}^{\top}\\ \vdots\\ \mathbf{X}_{b}^{\top}\end{bmatrix}\text{ \ and \ }\underset{(b-a)\times r}{\mathbf{F}_{a:b}}=\begin{bmatrix}\mathbf{f}_{a+1}^{\top}\\ \vdots\\ \mathbf{f}_{b}^{\top}\end{bmatrix}\text{ \ for all \ }0\leq a<b\leq T.$
It follows that
$\begin{bmatrix}\mathbf{X}_{0:k^{\ast}}\\ \mathbf{X}_{k^{\ast}:T}\end{bmatrix}=\begin{bmatrix}\mathbf{F}_{0:k^{\ast}}\bm{\Lambda}_{0}^{\top}\\ \mathbf{F}_{k^{\ast}:T}\bm{\Lambda}_{1}^{\top}\end{bmatrix}+\mathbf{E}=\begin{bmatrix}\mathbf{F}_{0:k^{\ast}}\mathbf{A}_{0}^{\top}\\ \mathbf{F}_{k^{\ast}:T}\mathbf{A}_{1}^{\top}\end{bmatrix}\bm{\Lambda}^{\top}+\mathbf{E}=:\mathbf{G}\bm{\Lambda}^{\top}+\mathbf{E},$
which is an observationally equivalent representation with a time-invariant loading matrix $\bm{\Lambda}$ and pseudo factors of dimension $r$ contained in $\mathbf{G}$.
We extend the observation made in Example 1 to the multiple change point setting and re-write the model in (1) as
$\mathbf{X}_{t}=\bm{\Lambda}\sum_{j=0}^{R}\mathbf{A}_{j}\mathbf{f}_{t}\cdot\mathbb{I}_{\{k_{j}<t\leq k_{j+1}\}}+\mathbf{e}_{t}=\bm{\Lambda}\mathbf{g}_{t}+\mathbf{e}_{t}$ (2)
with $\bm{\Lambda}_{j}=\bm{\Lambda}\mathbf{A}_{j}$ for all $0\leq j\leq R$. Under the homoscedasticity of $\{\mathbf{f}_{t}\}$, we have
$\mathsf{Cov}(\mathbf{g}_{k_{j}})=\mathbf{A}_{j-1}\mathsf{Cov}(\mathbf{f}_{0})\mathbf{A}_{j-1}^{\top}\neq\mathbf{A}_{j}\mathsf{Cov}(\mathbf{f}_{0})\mathbf{A}_{j}^{\top}=\mathsf{Cov}(\mathbf{g}_{k_{j}+1})$
for $1\leq j\leq R$. Then, the problem of detecting the multiple change points in the loadings under (2), becomes that of detecting change points in the covariance of the pseudo factors $\mathbf{g}_{t}$.
2.2 Methodology
Let us suppose that the number of pseudo factors $r$ is known. Then, we can estimate the pseudo factors $\mathbf{g}_{t}$ as, up to a rotation, $\sqrt{T}$ times the $r$ leading eigenvectors of the $T\times T$ matrix $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}=\mathbf{X}_{0:T}$. Denoting such estimator by $\widehat{\mathbf{g}}_{t}$, we propose to test for a change under the model in (1) and, if any, estimate the multiple change points, by scanning the data using the following MOSUM statistic:
$\mathcal{T}_{N,T,\gamma}(k)$ $=\left|\mathbf{M}_{N,T,\gamma}^{\top}(k)\mathbf{V}_{k}^{-1}\mathbf{M}_{N,T,\gamma}(k)\right|^{1/2}\text{ \ for \ }\gamma\leq k\leq T-\gamma,\text{ \ with}$ (3)
$\mathbf{M}_{N,T,\gamma}(k)$ $=\frac{1}{\sqrt{2\gamma}}\mathsf{Vech}\left(\sum_{t=k+1}^{k+\gamma}\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\sum_{t=k-\gamma+1}^{k}\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}\right),$
where $\gamma\geq 1$ is a pre-selected bandwidth. The matrix $\mathbf{V}_{k}$ denotes the long-run covariance matrix of $\mathbf{M}_{N,T,\gamma}(k)$ which, in the absence of any change point, satisfies $\mathbf{V}_{k}\equiv\mathbf{V}$ for all $k$, with $\mathbf{V}$ defined explicitly in Theorem 1 below.
For testing the null hypothesis of no change point, $\mathcal{H}_{0}:\,R=0$, we compare the maximum of the MOSUM statistics against some threshold, say $D_{T,\gamma}$, and reject $\mathcal{H}_{0}$ if
$\mathcal{T}_{N,T,\gamma}=\max_{\gamma\leq k\leq T-\gamma}\mathcal{T}_{N,T,\gamma}(k)>D_{T,\gamma}.$
In Theorem 1 below, we derive the asymptotic null distribution of $\mathcal{T}_{N,T,\gamma}$, which enables selecting $D_{T,\gamma}$ as its upper $\alpha$-quantile at a given significance level $\alpha\in(0,1)$.
Beyond testing for any change, we propose to detect and locate the multiple change points by adopting an approach proposed by Eichinger and Kirch, (2018) in the univariate mean change point setting. Simply put, we select every local maximiser of $\mathcal{T}_{N,T,\gamma}(k)$ over a sufficiently large enough interval at which $\mathcal{T}_{N,T,\gamma}(k)$ exceeds the threshold. Specifically, for some fixed $\eta\in(0,1]$, we set as a change point estimator every $\widehat{k}$ that simultaneously satisfies
$\widehat{k}={\arg\max}_{k:\,|\widehat{k}-k|\leq\eta\gamma}\,\mathcal{T}_{N,T,\gamma}(k)\text{ \ and \ }\mathcal{T}_{N,T,\gamma}(\widehat{k})>D_{T,\gamma}.$ (4)
[…]
For the implementation of the MOSUM procedure, we require an estimator of the long-run covariance matrix $\mathbf{V}_{k}$. While we allow it to be location-dependent to account for the heteroscedasticity in the presence of change points, the difficulty in estimating a long-run covariance matrix of multivariate time series is well-known. In the current setting, this is further augmented by that the computation of $\mathcal{T}_{N,T,\gamma}(k)$ calls for the inverse of $\mathbf{V}_{k}$, which may bring further numerical instabilities. Therefore, we opt to use the following HAC-type estimator in place of $\mathbf{V}_{k}$,
$\widehat{\mathbf{V}}$ $=\widehat{\bm{\Gamma}}(0)+\sum_{\ell=1}^{m}\left(1-\frac{\ell}{m+1}\right)\left(\widehat{\bm{\Gamma}}(\ell)+\widehat{\bm{\Gamma}}^{\top}(\ell)\right),\text{ \ where}$ (5)
$\widehat{\bm{\Gamma}}(\ell)$ $=\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{I}_{r}\right)\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t-\ell}\widehat{\mathbf{g}}_{t-\ell}^{\top}-\mathbf{I}_{r}\right)^{\top}$
with some bandwidth $m\geq 1$. We later show that the matrix $\widehat{\mathbf{V}}$ provides a consistent estimator of the long-run covariance matrix of $\mathbf{M}_{N,T,\gamma}(k)$ under $\mathcal{H}_{0}$ (see Proposition 2). We note that the use of the Bartlett kernel in (5) is only one of the many possible choices, and we refer to Bai et al., (2024) for a comprehensive analysis of various kernel-based estimators of $\mathbf{V}$. For consistency of multiple change point detection, we only require that a positive definite matrix is used in place of $\mathbf{V}_{k}$ (Theorem 3). By default, we propose to adopt $\widehat{\mathbf{V}}$ (or its diagonal entries) which is shown to work well; see Section 4 for further discussions.
3 Theoretical properties
3.1 Assumptions
We begin by providing a definition of $\mathcal{L}_{\nu}$-decomposable Bernoulli shifts.
Definition 1.
The $d$-dimensional sequence $\left\{\mathbf{m}_{t},-\infty<t<\infty\right\}$ forms an $L_{\nu}$-decomposable Bernoulli shift if and only if it holds that $\mathbf{m}_{t}=h(\bm{\eta}_{t},\bm{\eta}_{t-1},\ldots)$, where (i) $h:\,\mathbb{S}^{\infty}\rightarrow\mathbb{R}^{d}$ is a non random measurable function; (ii) $\{\bm{\eta}_{t}\}_{t\in\mathbb{Z}}$ is an i.i.d. sequence with values in a measurable space $\mathbb{S}$; (iii) $\mathsf{E}(\mathbf{m}_{t})=0$ and $|\|\mathbf{m}_{t}\||_{\nu}<\infty$; and (iv) $|\|\mathbf{m}_{t}-\mathbf{m}_{t,l}^{\ast}\||_{\nu}\leq c_{0}l^{-a}$ for some $c_{0}>0$ and $a>0$, where $\mathbf{m}_{t,l}^{\ast}=h(\bm{\eta}_{t},\ldots,\bm{\eta}_{t-l+1},\bm{\eta}_{t-l,t,l}^{\ast},\bm{\eta}_{t-l-1,t,l}^{\ast},\ldots)$, with $\{\bm{\eta}_{s,t,l}^{\ast},\,-\infty<s,l,t<\infty\}$ that are i.i.d. copies of $\bm{\eta}_{0}$ which are independent of $\{\bm{\eta}_{t}\}_{t\in\mathbb{Z}}$.
Since the seminal works by Wu, (2005) and Berkes et al., (2011) (see also Hörmann,, 2009), decomposable Bernoulli shifts have proven a very convenient way to model and study dependent time series, mainly due to their generality and since it is much easier to verify whether a sequence forms a decomposable Bernoulli shift than, e.g. verifying mixing conditions. Virtually all the most commonly employed models in econometrics and statistics can be shown to generate decomposable Bernoulli shifts, such as ARMA and (G)ARCH processes, non-linear time series models (e.g. random coefficient autoregressive models and threshold models), Volterra series and data generated by dynamical systems; see Berkes et al., (2011), Aue et al., (2009) and Liu and Lin, (2009).
We establish the theoretical properties of the proposed MOSUM procedure under the following assumptions.
Assumption 1.
  1. (i)
    There exists some fixed $\epsilon\in(0,1)$ such that $\{\mathbf{f}_{t}\}_{t\in\mathbb{Z}}$ is an $\mathcal{L}_{\nu}$-decomposable Bernoulli shift with $\nu\geq 8\rho+\epsilon$ for $\rho=1$ or $\rho=2$, and $a>2$.
  2. (ii)
    $\bm{\Sigma}_{F}=\mathsf{E}(\mathbf{f}_{t}\mathbf{f}_{t}^{\top})\in\mathbb{R}^{r\times r}$ is positive definite.
  3. (iii)
    Denoting the long-run variance matrix of $\mathcal{F}_{t}=\mathsf{Vech}(\mathbf{f}_{t}\mathbf{f}_{t}^{\top})$ by
    $\mathbf{D}=\lim_{T\rightarrow\infty}\frac{1}{T}\mathsf{E}\left\{\left[\sum_{t=1}^{T}(\mathcal{F}_{t}-\mathsf{E}(\mathcal{F}_{t}))\right]\;\left[\sum_{t=1}^{T}\left(\mathcal{F}_{t}-\mathsf{E}(\mathcal{F}_{t})\right)\right]^{\top}\right\},$ (6)
    we have $\mathbf{D}$ invertible.
Assumption 2.
There exists some $c_{0}\in(0,\infty)$ such that:
  1. (i)
    $\bm{\lambda}_{i}$ is deterministic with $\|\bm{\lambda}_{i}\|\leq c_{0}$ for all $1\leq i\leq N$ and $N\in\mathbb{N}$.
  2. (ii)
    $\|N^{-1}\bm{\Lambda}^{\top}\bm{\Lambda}-\bm{\Sigma}_{\Lambda}\|\leq c_{0}N^{-1/2}$ for all $N\in\mathbb{N}$, where $\bm{\Sigma}_{\Lambda}\in\mathbb{R}^{r\times r}$ is positive definite.
Assumption 3.
There exist some $\epsilon\in(0,1)$ and $c_{0}\in(0,\infty)$ such that the following holds for all $N,T\in\mathbb{N}$ and $\rho=1$ or $\rho=2$:
  1. (i)
    $\mathsf{E}(e_{it})=0$ and $\mathsf{E}(|e_{i,t}|^{8+\epsilon})<\infty$ for all $1\leq i\leq N$ and $1\leq t\leq T$.
  2. (ii)
    Letting $\gamma_{s,t}=N^{-1}\sum_{i=1}^{N}\mathsf{E}(e_{i,s}e_{i,t})$, it holds that $\sum_{t=1}^{T}|\gamma_{s,t}|\leq c_{0}$ for all $1\leq s\leq T$.
  3. (iii)
    $\mathsf{E}(|\sum_{i=1}^{N}(e_{i,t}e_{i,s}-\gamma_{s,t})|^{4+\epsilon})\leq c_{0}N^{2+\epsilon/2}$ for all $1\leq s,t\leq T$.
  4. (iv)
    $\mathsf{E}(\|\sum_{i=1}^{N}\bm{\lambda}_{i}e_{i,t}\|^{8+\epsilon})\leq c_{0}N^{4+\epsilon/2}$ for all $1\leq t\leq T$.
  5. (v)
    $\sum_{j=1}^{N}\left|\mathsf{E}\mathit{(e_{i,t}e_{j,t})}\right|\leq c_{0}$ for all $1\leq i\leq N$ and $1\leq t\leq T$.
  6. (vi)
    $\mathsf{E}(|\sum_{t=1}^{T}(e_{i,t}e_{j,t}-\mathsf{E}(e_{i,t}e_{j,t}))|^{4\rho+\epsilon})\leq c_{0}T^{2\rho+\epsilon/2}$ for all $1\leq i,j\leq N$.
Assumption 4.
There exist some $\epsilon\in(0,1)$ and $c_{0}\in(0,\infty)$ such that the following holds for all $N,T\in\mathbb{N}$ and $\rho=1$ or $\rho=2$:
  1. (i)
    $\mathsf{E}(\|\sum_{t=a+1}^{b}\mathbf{g}_{t}\sum_{i=1}^{N}(e_{i,t}e_{i,s}-\gamma_{s,t})\|^{2\rho+\epsilon})\leq c_{0}(N(b-a))^{\rho+\epsilon/2}$ for all $1\leq s\leq T$ and $0\leq a<b\leq T$.
  2. (ii)
    $\mathsf{E}(|\sum_{i=1}^{N}\sum_{t=a+1}^{b}\bm{\lambda}_{i}^{\top}\mathbf{g}_{t}e_{i,t}|^{4\rho+\epsilon})\leq c_{0}(N(b-a))^{2\rho+\epsilon/2}$ for all $0\leq a<b\leq T$.
  3. (iii)
    $\mathsf{E}(|\sum_{t=1}^{T}\bm{\lambda}_{i}^{\top}\mathbf{g}_{t}e_{j,t}|^{4\rho+\epsilon})\leq c_{0}T^{2\rho+\epsilon/2}$ for all $1\leq i,j\leq N$.
Assumptions 14 are closely related to those adopted in the factor model literature, see also Bai, (2003). Assumption 2 is the same as Assumptions B in Bai, (2003). In Assumption 1 (i), we strengthen the moment condition typically employed in the literature on $\mathbf{f}_{t}$, switching from the existence of the $4$-th moment to that of the $8$-th or higher moment. Also, we assume a specific form of dependence for $\{\mathbf{f}_{t},-\infty<t<\infty\}$ which, as mentioned above, accommodates a wide range of time series models. This is required to derive a Strong Invariance Principle (SIP) for $\mathsf{Vech}(\mathbf{f}_{t}\mathbf{f}_{t}^{\top})$, see Lemma A.9 in Appendix A.1. Assumption 3 allows for weak temporal and cross-sectional dependence in the idiosyncratic component, with similarities between (ii) and Assumption E1 in Bai, (2003), (iii) and C5, and (v) and E2; part (iv) strengthens their F3 and also Assumption 6 (ii) of Bai et al., (2024); part (vi) can be derived under more primitive conditions on $e_{i,t}$. Assumption 4 (i)(ii) extend Assumption F1 and F2 of Bai, (2003), to account for the scanning for multiple change points performed by the MOSUM procedure.
Generally, the strengthened conditions found in Assumptions 1 (i) and 3 on the moments of $\mathbf{f}_{t}$ and $e_{i,t}$, are required as we go a step further from the typical factor modelling literature that focuses on establishing the consistency of the estimated factors, to control the partial sums involved in the MOSUM process. We note that $\rho=1$ in Assumptions 1, 3 and 4, is sufficient for deriving the asymptotic null distribution of the MOSUM test statistic (Theorem 1) as well as the detection consistency of the MOSUM procedure (Theorem 3 (a)), while $\rho=2$ is required for establishing the rate of estimation for the change points (Theorem 3 (b)).
Assumption 5.
  1. (i)
[…]
  2. (ii)
    $\|\mathbf{A}_{j}\|\leq c_{0}\in(0,\infty)$ for all $0\leq j\leq R$.
  3. (iii)
    Denoting by $\bm{\Sigma}_{G}=T^{-1}\sum_{t=1}^{T}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})$, the eigenvalues of $\bm{\Sigma}_{G}\bm{\Sigma}_{\Lambda}$ are positive and distinct.
Assumption 6.
There exists some $\epsilon_{\circ}\in(0,\infty)$ such that
$\lim_{\min(N,T)\rightarrow\infty}\frac{T^{1/2+\epsilon_{\circ}}}{N}=0.$
When $R=0$, Assumption 5 only requires that $\bm{\Sigma}_{F}\bm{\Sigma}_{\Lambda}$ has distinct eigenvalues, paralleling the commonly found condition such as Assumption G of Bai, (2003). When $R\geq 1$, part (i) assumes that the change points are linearly spaced. The positive definiteness imposed on $\bm{\Sigma}_{G}$ in part (iii), together with part (i) and Assumption 1 (ii), implies that any local factors are pervasive over segment(s) where they are present. Finally, Assumption 6 is also found in Bai et al., (2024), and arises from that we construct the MOSUM process based on an estimate of the latent factors.
3.2 Asymptotic null distribution
We present the limiting distribution of the maximally selected MOSUM statistic in (3). We write, for simplicity, $d=r(r+1)/2$ and denote by $\beta=\log(N)/\log(T)$; under Assumption 6, we have $1/2+\epsilon_{\circ}<\beta$ with some $\epsilon_{\circ}>0$. Define also
$\zeta=\max\left\{\frac{2}{\nu},1-\min(1,\beta)\right\},$ (7)
where $\nu$ is defined in Assumption 1. Then, we always have $\zeta\in(0,1/2)$.
Theorem 1.
Suppose that Assumption 16 hold with $\rho=1$ for Assumptions 1, 3 and 4, and let the bandwidth $\gamma$ satisfy
${\frac{T^{2\zeta}\log(T/\gamma)}{\gamma}\rightarrow 0\text{ \ and \ }\frac{\gamma}{T}\rightarrow 0.}$ (8)
Let us define $\mathbf{V}=\mathbf{L}_{r}(\mathbf{H}_{0}^{\top}\otimes\mathbf{H}_{0}^{\top})\mathbf{K}_{r}\mathbf{D}\mathbf{K}_{r}^{\top}(\mathbf{H}_{0}\otimes\mathbf{H}_{0})\mathbf{L}_{r}^{\top}$, with $\mathbf{D}$ in (6), and
$\mathbf{H}_{0}=\quad\text{p}\!\!\!\!\!\!\!\!\!\!\!\lim_{\min(N,T)\rightarrow\infty}\frac{1}{NT}(\bm{\Lambda}^{\top}\bm{\Lambda})(\mathbf{G}^{\top}\widehat{\mathbf{G}})\bm{\Phi}_{NT}^{-1}$ (9)
where we denote with $\bm{\Phi}_{NT}\in\mathbb{R}^{r\times r}$ the diagonal matrix containing the $r$ largest eigenvalues of $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ on its diagonal.
  1. (a)
    Under $\mathcal{H}_{0}:\,R=0$, for all $x\in\mathbb{R}$, we have
    $\lim_{\min(N,T)\rightarrow\infty}\mathsf{P}\left(a\left(\frac{T}{\gamma}\right)\max_{\gamma\leq k\leq T-\gamma}\left|\mathbf{M}_{N,T,\gamma}^{\top}(k)\mathbf{V}^{-1}\mathbf{M}_{N,T,\gamma}(k)\right|^{1/2}-b_{d}\left(\frac{T}{\gamma}\right)\leq x\right)\\ =\exp\left(-2\exp(-x)\right),$ (10)
    where $a(x)=\sqrt{2\log(x)}$ and $b_{d}(x)=2\log(x)+d\log\log(x)/2+\log(1/2)-\log(\Gamma(d/2))$.
  2. (b)
    The assertion in (a) continues to hold if $\mathbf{V}$ is replaced by a positive definite matrix $\widehat{\mathbf{V}}$ satisfying
    $\left\|\widehat{\mathbf{V}}-{\mathbf{V}}\right\|=o_{P}\left(\log^{-1}(T/\gamma)\right).$ (11)
The limiting law in Theorem 1 is analogous to those derived in Hušková and Slabỳ, (2001) and Eichinger and Kirch, (2018), modulo the fact that here, we deal with $d$-variate vectors and therefore the function $b_{d}(\cdot)$ depends on $d$. In contrast with a “standard” multivariate time series application, however, in our result, the cross-sectional dimension $N$ also plays a role through the definition of $\zeta$ in (7). As a by-product, we establish the SIP of the process $\mathsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ after an appropriate centering (see Lemma A.11 in Appendix A.1), which proves crucial in deriving the asymptotic null distribution in (10).
Based on this limiting law of the maximally selected MOSUM process, we reject $\mathcal{H}_{0}:\,R=0$ at the significance level $\alpha\in(0,1)$, if
$\max_{\gamma\leq k\leq T-\gamma}\left|\mathbf{M}_{N,T,\gamma}^{\top}(k)\mathbf{V}^{-1}\mathbf{M}_{N,T,\gamma}(k)\right|^{1/2}>\widetilde{D}_{T,\gamma}(\alpha):=\frac{b_{d}(T/\gamma)-\log\log\left(\frac{1}{\sqrt{1-\alpha}}\right)}{a(T/\gamma)}.$ (12)
The condition in (8) places both upper and lower bounds on the bandwidth $\gamma$. Specifically, $\gamma$ is required to grow sufficiently faster than $T^{2\zeta}$ while satisfying $T^{-1}\gamma\rightarrow 0$, and the former restriction calls for larger $\gamma$ when $\mathbf{f}_{t}$ has fewer moments or when $N$ is small. We note that $\beta$ is known and does not need to be estimated. Conversely, $\nu$ is in general not known. We may select $\gamma$ satisfying (8) by plugging in an estimate of $\nu$, say $\widehat{\nu}$; alternatively, one can decide a value of $\nu$, say $\nu^{\ast}$, and test whether $\left|\mathbf{g}_{t}\right|_{\nu^{\ast}}<\infty$. In both cases, one difficulty is that $\mathbf{g}_{t}$ is not observable; however, deriving $\nu$ from the data $\mathbf{X}_{t}$ yields a lower bound.
Theorem 1 (b) shows that when the unknown $\mathbf{V}$ is replaced by its estimator, the asymptotic null distribution continues to hold provided that (11) is met. The following Proposition 2 guarantees that this requirement is met by the estimator $\widehat{\mathbf{V}}$ proposed in (5), strengthening the observation made in Bai et al., (2024) that $\|\widehat{\mathbf{V}}-\mathbf{V}\|=o_{P}(1)$.
Proposition 2.
Suppose that Assumption 16 hold with $\rho=1$ for Assumptions 1, 3 and 4. Also, let the bandwidth $m$ satisfy
$\frac{\log(T/\gamma)}{m}\rightarrow 0\text{ \ and \ }\frac{m\log(T/\gamma)}{\sqrt{\min(N,T)}}\rightarrow 0.$ (13)
Then, as $\min(N,T)\rightarrow\infty$, the estimator $\widehat{\mathbf{V}}$ in (5) satisfies the condition in (11).
3.3 Consistency in multiple change point estimation
To establish the consistency of the MOSUM procedure in multiple change point detection, we make the following assumption on the size of changes.
Assumption 7.
  1. (i)
    $\min_{0\leq j\leq R}\Delta_{j}\geq 2\gamma$, where $\Delta_{j}=k_{j+1}-k_{j}$.
  2. (ii)
    At each $1\leq j\leq R$, let $\bm{\delta}_{j}=\mathbf{A}_{j}\bm{\Sigma}_{F}\mathbf{A}_{j}^{\top}-\mathbf{A}_{j-1}\bm{\Sigma}_{F}\mathbf{A}_{j-1}^{\top}$ and $d_{j}=\|\bm{\delta}_{j}\|$. Then for $\omega_{T}^{(1)}\rightarrow\infty$ arbitrarily slowly, it holds that
    $\frac{\omega_{T}^{(1)}\sqrt{\log(T/\gamma)}}{\min_{1\leq j\leq R}d_{j}\sqrt{\gamma}}=o(1).$
Assumption 7 (i) ensures that there exists at most a single change point over each moving window, and is implied jointly by Assumption 5 (i) and the condition (8) on $\gamma$. Condition (ii) permits local changes with $d_{j}\rightarrow 0$, at a sufficiently slow rate, which is the case e.g. when the neighbouring loading matrices are rotations of one another with $\|\mathbf{A}_{j}-\mathbf{A}_{j-1}\|\to 0$.
Theorem 3.
Suppose that Assumption 16 hold with $\rho=1$ for Assumptions 1, 3 and 4. Additionally, let Assumption 7 hold and the bandwidth $\gamma$ satisfy (8), and suppose that some positive definite matrix $\widetilde{\mathbf{V}}$ is used in place of $\mathbf{V}_{k}$ in $\mathcal{T}_{N,T,\gamma}(k)$, see (3).
  1. (a)
[…]
    $\mathsf{P}\left(\widehat{R}=R;\,\max_{1\leq j\leq R}|\widehat{k}_{j}-k_{j}|\leq\eta\gamma\right)\rightarrow 1\text{ \ as \ }\min(N,T)\rightarrow\infty.$
  2. (b)
    Further, if Assumptions 1, 3 and 4 hold with $\rho=2$, there exists some sequence $\omega_{T}^{(2)}\rightarrow\infty$ arbitrarily slowly, such that
    $\mathsf{P}\left(\widehat{R}=R;\,\max_{1\leq j\leq R}d_{j}^{2}|\widehat{k}_{j}-k_{j}|\leq\omega_{T}^{(2)}\right)\rightarrow 1\text{ \ as \ }\min(N,T)\rightarrow\infty.$
Theorem 3 shows that the MOSUM procedure consistently estimates the total number and the locations of the change points. Here, we adopt a fixed, positive definite matrix $\widetilde{\mathbf{V}}$ in place of $\mathbf{V}_{k}$ which, without being a consistent estimator of the latter at some $k$, still leads to consistency in multiple change point detection. This flexibility in the choice of $\widetilde{\mathbf{V}}$ is particularly favourable since, as noted earlier in Section 2.2, the estimation of (time-varying) long-run covariance matrix for multivariate time series is challenging. At the same time, it incurs an additional multiplicative factor of $\omega_{T}^{(1)}$ in the detection lower bound placed on $d_{j}\sqrt{\gamma}$, see Assumption 7 (ii). Under a stronger moment assumption, we obtain the rate of estimation which is inversely proportional to the squared size of change as $|\widehat{k}_{j}-k_{j}|=O_{P}(d_{j}^{-2})$, This indicates that dominant changes are located with better accuracy, such as those accompanied by a change in the dimension of the factor space due to the introduction or disappearance of factors(s).
4 Numerical experiments
4.1 Tuning parameter selection
Empirical performance of the MOSUM procedure depends on the choice of tuning parameters.
Inspecting the proof of Theorem 1, we observe that the requirement on $\nu$ in Assumption 1 (i) may be weakened if the $r$ largest eigenvalues of $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ are bounded away from zero deterministically. Inspired by this and the condition in (8), we propose to select the bandwidth as $\gamma=\lfloor T^{2\zeta}\cdot\log^{\varrho}(T)\rfloor$ with $\zeta=\max(2/5,1-\log(N)/\log(T))$. Thus-selected bandwidth with $\varrho=1.1$ works reasonably well in our simulation studies where datasets of dimensions $N\leq 500$ and $T\leq 1000$ are considered. For the real data application in Section 5 with $T\geq 4000$, we set $\varrho=0.5$.
We set the threshold $D_{T,\gamma}$ as described in Theorem 3, namely $D_{T,\gamma}=\widetilde{D}_{T,\gamma}(\alpha)\cdot\omega_{T}^{(1)}$ where $\widetilde{D}_{T,\gamma}(\alpha)$ is given by (12) according to the asymptotic null distribution in Theorem 1. For $\omega_{T}^{(1)}$, we have considered $\log^{\kappa}(T/\gamma)$ with $\kappa\in\{0,0.1,0.2,0.3\}$, and observed that the choice of $\kappa=0.2$ returned stably good performance in all experiments. Due to the presence of this $\omega_{T}^{(1)}$, the selection of the significance level $\alpha\in(0,1)$ has relatively little influence and in all our studies, we set $\alpha=0.05$. Finally, for the detection rule in (4), we set $\eta=0.6$.
For estimating the number of pseudo factors $r$, we apply the approach proposed by Alessi et al., (2010) in combination with the three information criteria of Bai and Ng, (2002): Addressing the arbitrariness in the choice of the penalty, it looks for a stable estimate of the factor number as the minimiser of the information criterion over sub-samples of varying dimensions and sample sizes. We take the median of the estimates from the three information criteria if they do not agree. On simulated datasets, we find that this approach consistently identifies the correct number of factors over $90\%$ of the realisations.
Finally, in place of $\mathbf{V}_{k}$ in (3), we plug in the estimator $\widehat{\mathbf{V}}$ in (5) with bandwidth $m=\lfloor T^{1/4}\rfloor$. Due to the presence of change points, the number of pseudo factors $r$ can be large in which case inverting the $d\times d$-matrix $\widehat{\mathbf{V}}$ with $d=r(r+1)/2$, may bring numerical instability. Therefore, we explore two approaches, one performing the standardisation using the full matrix $\widehat{\mathbf{V}}$, and the other using its diagonal entries only, respectively referred to as MOSUM-full and MOSUM-diagonal in simulation studies. Note that MOSUM-diagonal meets the requirement in Theorem 3 provided that all diagonal entries of $\widehat{\mathbf{V}}$ are positive.
4.2 Settings
We consider the following data generating processes considered in Li et al., (2023) and Duan et al., (2023).
  1. (M1)
[…]
    $\chi_{it}$ $=\bm{\Lambda}_{j}\mathbf{f}_{t}\text{ \ with \ }\mathbf{f}_{t}\sim\mathcal{N}_{r_{0}}(\mathbf{0},\bm{\Sigma}_{j})\text{ \ for \ }k_{j-1}+1\leq t\leq k_{j},\text{ \ and}$
    $\bm{\Sigma}_{0}$ $=[\sigma_{0,ij}]=\mathbf{D}\bm{\Sigma}_{F}\mathbf{D}\text{ \ with\ }\mathbf{D}=\text{diag}(d_{ii},\,1\leq i\leq r_{0}),\,d_{ii}\sim_{\text{iid}}\mathcal{U}[0.5,1.5],$
    $\bm{\Sigma}_{1}=\bm{\Sigma}_{2}$ $=[\sigma_{1,ij}]\text{ \ with \ }\sigma_{1,ij}=\sigma_{1,ji}=\left\{\begin{array}[]{ll}0.9\sqrt{\sigma_{0,11}\sigma_{0,22}}&\text{for \ }(i,j)=(1,2),\\ 1.3^{2}\sigma_{0,55}&\text{for \ }(i,j)=(5,5),\\ 0.5^{|i-5|}\sqrt{\sigma_{0,ii}\sigma_{0,55}}&\text{for \ }1\leq i\leq 4,\\ \sigma_{0,ij}&\text{otherwise.}\end{array}\right.$
[…]
  2. (M2)
    We join together three single change point scenarios from Duan et al., (2023) to form a multiple change point one: Setting $R=3$ and $r_{0}=3$, we generate
    $\mathbf{f}_{t}$ $=\rho_{f}\mathbf{f}_{t-1}+\bm{\varepsilon}_{f,t},\,\bm{\varepsilon}_{f,t}\sim_{\text{iid}}\mathcal{N}_{r_{0}}(\mathbf{0},\mathbf{I}_{r_{0}}),$
    $\mathbf{e}_{t}$ $=\rho_{e}\mathbf{e}_{t-1}+\bm{\varepsilon}_{e,t},\,\bm{\varepsilon}_{e,t}\sim\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma}_{e})\text{ \ with \ }\bm{\Sigma}_{e}=[(0.3)^{|i-j|},\,1\leq i,j\leq p],$
[…]
    $\mathbf{C}_{1}=\begin{bmatrix}0.5&0&0\\ c_{1,21}&1&0\\ c_{1,31}&c_{1,32}&1.5\end{bmatrix},\quad\mathbf{C}_{2}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix}\text{ \ and \ }\mathbf{C}_{3}=[c_{3,ij},\,1\leq i,j\leq r_{0}],$
    with $c_{1,ij}\sim_{\text{iid}}\mathcal{N}(0,1)$ and $c_{3,ij}\sim_{\text{iid}}\mathcal{N}(0,1/r_{0})$. The factor number varies as $(r_{0},\ldots,r_{3})=(3,3,2,3)$, while the number of pseudo factors increases from $3$ to $6$ due to the change at $k_{3}$. We vary $T\in\{400,600,800,1000\}$ and $N\in\{100,200,500\}$ as well as $(\rho_{f},\rho_{e})=\{(0,0),(0.7,0.3)\}$.
  3. (M3)
    Additionally, we consider the “null” model with $R=0$ by generating the data from the model corresponding to the first segment of (M2) for each setting.
4.3 Results
For each setting, we generate $200$ realisations and report the distribution of $\widehat{R}-R$, and the accuracy of change point estimators measured by
$\frac{1}{200}\sum_{i=1}^{200}\mathbb{I}\left[\min_{1\leq\ell\leq\widehat{K}^{(i)}}|\widehat{k}_{\ell}^{(i)}-k_{j}|\leq\log(T)\right]$
[…]
Tables 14 report the summary of the results over $200$ realisations. Overall we observe that the MOSUM procedure (‘MOSUM’) performs competitively against BSCOV under (M1), and the former outperforms the latter in the other scenarios. In the presence of change points under (M2), we observe that BSCOV tends to over-estimate the number of change points particularly as $T$ increases, and this tendency becomes greater when the data are serially dependent (with $(\rho_{f},\rho_{e})=(0.7,0.3)$). In the absence of change points under (M3), BSCOV detects spurious change points when $T$ is smaller. On the contrary, MOSUM correctly estimates the number of change points on most realisations even when the data are serially dependent, regardless of whether $R=0$ or not. Additionally, we observe that $k_{j}$’s which are associated with changes in the number of pseudo factors (such as $k_{2}$ under (M1) and $k_{3}$ under (M2)) are estimated with higher accuracy, which agrees with the observations made in Duan et al., (2023) and also below Theorem 3.
Between MOSUM-full and MOSUM-diagonal, the latter demonstrates better accuracy in estimating both the total number and locations of change points when $R\geq 1$. This is explained by that $d=r(r+1)/2$ is as large as $d=28$ under (M1) and $d=21$ under (M2). Retaining the diagonal elements of $\widehat{\mathbf{V}}$ only, effectively performs standardisation without suffering from the numerical instability inherent in inverting a large matrix. When $R=0$, MOSUM-full performs marginally better as here, the factor number is kept at $r=3$ under (M3), leading to $d=6$.
Table 1: (M1) with $R=2$: Summary of change point estimators returned by MOSUM and BSCOV. The results for BSCOV have been taken from Li et al., (2023).
$\widehat{R}-R$ Accuracy
Method LRV $-2\leq$ $-1$ 00 $1$ $\geq 2$ $j=1$ $j=2$
MOSUM Diagonal 0 0 0.985 0.015 0 0.7 0.925
Full 0 0.005 0.94 0.055 0 0.595 0.8
BSCOV 0 0.03 0.97 0 0 0.64 0.95
Table 2: (M2) with $(\rho_{f},\rho_{e})=(0,0)$ and $R=3$: Summary of change point estimators returned by MOSUM and BSCOV over $200$ realisations.
$\widehat{R}-R$ Accuracy
$T$ $N$ Method LRV $-2\leq$ $-1$ 00 $1$ $\geq 2$ $j=1$ $j=2$ $j=3$
$400$ $100$ MOSUM Diagonal 0 0.01 0.99 0 0 0.82 0.88 0.985
Full 0 0.025 0.975 0 0 0.75 0.895 0.93
BSCOV 0.02 0 0.98 0 0 0.725 0.85 0.98
$200$ MOSUM Diagonal 0 0 1 0 0 0.795 0.875 0.97
Full 0 0.035 0.965 0 0 0.73 0.86 0.93
BSCOV 0.01 0.005 0.985 0 0 0.79 0.855 0.99
$500$ MOSUM Diagonal 0 0.02 0.98 0 0 0.8 0.88 0.975
Full 0 0.03 0.97 0 0 0.74 0.895 0.95
BSCOV 0.01 0 0.98 0.01 0 0.8 0.89 0.98
$600$ $100$ MOSUM Diagonal 0 0 1 0 0 0.83 0.88 0.99
Full 0 0 0.995 0.005 0 0.75 0.87 0.945
BSCOV 0 0 0.99 0.01 0 0.81 0.895 0.995
$200$ MOSUM Diagonal 0 0 1 0 0 0.81 0.89 0.98
Full 0 0.005 0.99 0.005 0 0.715 0.905 0.95
BSCOV 0 0 0.995 0.005 0 0.875 0.945 0.995
$500$ MOSUM Diagonal 0 0 1 0 0 0.83 0.92 0.99
Full 0 0 1 0 0 0.755 0.885 0.945
BSCOV 0.005 0 0.995 0 0 0.865 0.95 0.995
$800$ $100$ MOSUM Diagonal 0 0 0.995 0.005 0 0.785 0.905 0.975
Full 0 0 0.985 0.01 0.005 0.715 0.88 0.925
BSCOV 0 0.015 0.795 0.185 0.005 0.885 0.91 0.995
$200$ MOSUM Diagonal 0 0 1 0 0 0.805 0.92 0.965
Full 0 0 0.985 0.015 0 0.705 0.875 0.9
BSCOV 0 0 0.78 0.215 0.005 0.89 0.965 0.995
$500$ MOSUM Diagonal 0 0 1 0 0 0.775 0.89 0.975
Full 0 0 0.98 0.02 0 0.765 0.92 0.92
BSCOV 0 0 0.8 0.2 0 0.86 0.95 0.995
$1000$ $100$ MOSUM Diagonal 0 0 1 0 0 0.865 0.9 0.97
Full 0 0 0.94 0.06 0 0.715 0.88 0.925
BSCOV 0 0.005 0.835 0.16 0 0.87 0.925 0.99
$200$ MOSUM Diagonal 0 0 0.985 0.015 0 0.83 0.935 0.98
Full 0 0 0.925 0.075 0 0.735 0.88 0.92
BSCOV 0 0 0.86 0.13 0.01 0.91 0.955 0.99
$500$ MOSUM Diagonal 0 0 0.995 0.005 0 0.805 0.895 0.98
Full 0 0 0.925 0.075 0 0.73 0.905 0.935
BSCOV 0 0.005 0.83 0.16 0.005 0.875 0.965 0.99
Table 3: (M2) with $(\rho_{f},\rho_{e})=(0.7,0.3)$ and $R=3$: Summary of change point estimators returned by MOSUM and BSCOV over $200$ realisations.
$\widehat{R}-R$ Accuracy
$T$ $N$ Method LRV $-2\leq$ $-1$ 00 $1$ $\geq 2$ $j=1$ $j=2$ $j=3$
$400$ $100$ MOSUM Diagonal 0.005 0.08 0.915 0 0 0.595 0.765 0.905
Full 0 0.06 0.935 0.005 0 0.49 0.74 0.85
BSCOV 0.005 0.12 0.84 0.035 0 0.6 0.375 0.92
$200$ MOSUM Diagonal 0.005 0.11 0.885 0 0 0.605 0.73 0.93
Full 0 0.125 0.875 0 0 0.44 0.67 0.88
BSCOV 0 0.145 0.825 0.03 0 0.625 0.32 0.9
$500$ MOSUM Diagonal 0 0.115 0.885 0 0 0.535 0.71 0.95
Full 0.005 0.115 0.88 0 0 0.485 0.68 0.855
BSCOV 0 0.11 0.85 0.04 0 0.605 0.32 0.915
$600$ $100$ MOSUM Diagonal 0 0.025 0.97 0.005 0 0.565 0.69 0.905
Full 0 0.045 0.925 0.03 0 0.48 0.705 0.85
BSCOV 0.005 0.095 0.75 0.14 0.01 0.695 0.355 0.95
$200$ MOSUM Diagonal 0 0.03 0.96 0.01 0 0.535 0.75 0.935
Full 0 0.025 0.945 0.03 0 0.535 0.68 0.875
BSCOV 0 0.1 0.785 0.11 0.005 0.665 0.38 0.945
$500$ MOSUM Diagonal 0 0.005 0.995 0 0 0.57 0.77 0.935
Full 0 0.025 0.945 0.03 0 0.495 0.685 0.865
BSCOV 0.02 0 0.825 0.145 0.01 0.6 0.735 0.96
$800$ $100$ MOSUM Diagonal 0 0.01 0.96 0.03 0 0.54 0.69 0.92
Full 0 0.005 0.885 0.11 0 0.49 0.685 0.85
BSCOV 0.01 0 0.115 0.625 0.25 0.61 0.705 0.97
$200$ MOSUM Diagonal 0 0 0.955 0.045 0 0.575 0.785 0.925
Full 0 0.03 0.85 0.12 0 0.455 0.69 0.83
BSCOV 0.01 0 0.08 0.595 0.315 0.625 0.805 0.975
$500$ MOSUM Diagonal 0 0.01 0.95 0.04 0 0.6 0.725 0.915
Full 0 0.01 0.835 0.155 0 0.495 0.63 0.8
BSCOV 0.015 0 0.085 0.58 0.32 0.655 0.775 0.97
$1000$ $100$ MOSUM Diagonal 0 0 0.945 0.055 0 0.585 0.765 0.91
Full 0 0.005 0.805 0.18 0.01 0.49 0.725 0.81
BSCOV 0.005 0.005 0.195 0.5 0.295 0.64 0.825 0.95
$200$ MOSUM Diagonal 0 0 0.945 0.05 0.005 0.545 0.77 0.92
Full 0 0.01 0.81 0.17 0.01 0.47 0.74 0.84
BSCOV 0 0.02 0.21 0.405 0.365 0.64 0.825 0.945
$500$ MOSUM Diagonal 0 0 0.93 0.07 0 0.605 0.73 0.915
Full 0 0 0.78 0.21 0.01 0.515 0.68 0.84
BSCOV 0 0.01 0.205 0.45 0.335 0.67 0.82 0.96
Table 4: (M3) with $R=0$: Distribution of $\widehat{R}-R$ returned by MOSUM and BSCOV over $200$ realisations for $(\rho_{f},\rho_{e})\in\{(0,0),(0.7,0.3)\}$.
$(\rho_{f},\rho_{e})=(0,0)$ $(\rho_{f},\rho_{e})=(0.7,0.3)$
$\widehat{R}-R$ $\widehat{R}-R$
$T$ $N$ Method LRV 00 $1$ $\geq 2$ 00 $1$ $\geq 2$
$400$ $100$ MOSUM Diagonal 0.985 0.015 0 0.895 0.08 0.025
Full 0.99 0.01 0 0.965 0.025 0.01
BSCOV 1 0 0 0.75 0.215 0.035
$200$ MOSUM Diagonal 0.995 0.005 0 0.88 0.11 0.01
Full 0.995 0.005 0 0.955 0.04 0.005
BSCOV 0.995 0.005 0 0.705 0.25 0.045
$500$ MOSUM Diagonal 0.99 0.01 0 0.88 0.11 0.01
Full 0.99 0.01 0 0.95 0.045 0.005
BSCOV 0.995 0.005 0 0.75 0.22 0.03
$600$ $100$ MOSUM Diagonal 0.99 0.01 0 0.855 0.125 0.02
Full 0.99 0.01 0 0.905 0.075 0.02
BSCOV 1 0 0 0.82 0.155 0.025
$200$ MOSUM Diagonal 0.99 0.01 0 0.845 0.14 0.015
Full 0.995 0.005 0 0.88 0.105 0.015
BSCOV 1 0 0 0.77 0.18 0.05
$500$ MOSUM Diagonal 0.985 0.015 0 0.855 0.11 0.035
Full 0.995 0.005 0 0.875 0.1 0.025
BSCOV 1 0 0 0.85 0.135 0.015
$800$ $100$ MOSUM Diagonal 1 0 0 0.905 0.085 0.01
Full 1 0 0 0.95 0.045 0.005
BSCOV 1 0 0 0.835 0.15 0.015
$200$ MOSUM Diagonal 1 0 0 0.94 0.05 0.01
Full 0.995 0.005 0 0.97 0.02 0.01
BSCOV 1 0 0 0.875 0.115 0.01
$500$ MOSUM Diagonal 1 0 0 0.92 0.075 0.005
Full 1 0 0 0.94 0.055 0.005
BSCOV 1 0 0 0.86 0.12 0.02
$1000$ $100$ MOSUM Diagonal 1 0 0 0.895 0.1 0.005
Full 1 0 0 0.945 0.05 0.005
BSCOV 1 0 0 0.915 0.085 0
$200$ MOSUM Diagonal 1 0 0 0.9 0.09 0.01
Full 1 0 0 0.955 0.045 0
BSCOV 1 0 0 0.895 0.09 0.015
$500$ MOSUM Diagonal 1 0 0 0.91 0.085 0.005
Full 1 0 0 0.95 0.05 0
BSCOV 1 0 0 0.91 0.085 0.005
5 Real data application
We consider daily stock prices from $72$ US blue chip companies across industry sectors between January 3, 2005 and February 16, 2022, retrieved from the Wharton Research Data Services. We measure the volatility as the daily high-low range as $\sigma_{it}^{2}=0.361(p_{it}^{\text{high}}-p_{it}^{\text{low}})^{2}$ where $p_{it}^{\text{high}}$ (resp. $p_{it}^{\text{low}}$) denotes the maximum (resp. minimum) price of stock $i$ on day $t$, and set $X_{it}=\log(\sigma_{it}^{2})$, see, e.g. Diebold and Yılmaz, (2014).
The panel data is unbalanced with the dimension $N=72$ being considerably smaller than the sample size $T=4312$, a situation that does not favour the use of the sub-sampling-based factor number estimator as pointed out by Onatski, (2024). Indeed, it returns $r=7$ from the dataset which may be an over-estimate. An alternative estimator based on taking the ratio of successive eigenvalues of $(NT)^{-1}(\mathbf{X}\mathbf{X}^{\top})$ returns $r=1$ (Ahn and Horenstein,, 2013), which implies that any change point we detect is solely due to heteroscedasticity of the factor. Therefore, we apply the proposed MOSUM procedure with varying $r\in\{1,\ldots,7\}$ while setting other tuning parameters as described in Section 4.1; this gives a bandwidth $\gamma=227$ corresponding almost to one trading year.
Figure 1 illustrates the series of standardised MOSUM statistics $D_{T,\gamma}^{-1}\cdot\mathcal{T}_{N,T,\gamma}(k),\,\gamma\leq k\leq T-\gamma$, obtained for different values of $r$, where the standardisation is applied to ensure that the MOSUM series derived from $\textsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ of different dimensions are comparable. See also Table 5 for the list of change point estimators. We observe that with increasing $r$, the procedure detects more change points. Further, denoting by $\widehat{\mathcal{K}}(r)$ the set of change point estimators detected with the factor number $r$, generally it holds that $\widehat{\mathcal{K}}(r)\subset\widehat{\mathcal{K}}(r^{\prime})$ for any $r<r^{\prime}$, when accommodating the possible bias in the change point estimators (up to $3$ months). There are prominent estimators in mid-2008 and 2009 detected invariably with all $r$, which are associated with the Great Financial Crisis in 2007–2009. With $r\geq 5$, 2020-02-24 and 2021-01-19 emerge as change points, which are accounted for by the stock market crash in February 2020 and the ensuing recession due to the COVID-19 pandemic.
Figure 1: US blue chip data: Standardised MOSUM statistics obtained with $r\in\{1,\ldots,7\}$ (top to bottom). Vertical lines denote the change point estimators and the horizontal line is at $y=1$.
Table 5: US blue chip data: List of change point estimators obtained with $r\in\{1,\ldots,7\}$.
$r=1$ 2008-07-07, 2009-06-05
$r=2$ 2008-07-03, 2009-06-08
$r=3$ 2008-07-03, 2009-06-08
$r=4$ 2008-07-03, 2009-06-04, 2012-10-05
$r=5$ 2008-07-03, 2009-06-04, 2012-07-13, 2016-05-20, 2017-06-08, 2020-02-24, 2021-01-19
$r=6$ 2006-02-01, 2007-03-09, 2008-07-03, 2009-08-07, 2011-11-10, 2012-10-05, 2014-08-19
2016-05-20, 2017-06-08, 2020-02-24, 2021-01-19
$r=7$ 2006-02-01, 2007-03-09, 2008-09-12, 2009-05-28, 2011-11-10, 2012-10-05, 2014-08-19
2016-02-17, 2017-06-08, 2020-02-24, 2021-01-19
6 Conclusions
The paper proposes a MOSUM procedure for change point analysis under a static factor model that is popularly adopted in econometrics and statistics. In addition to deriving the asymptotic null distribution of the maximally selected MOSUM statistic, we establish the consistency of the procedure in multiple change point estimation with the accompanying rate of estimation, contributing to the relatively scarce literature on multiple change point detection in factor models. On a range of simulated datasets and in a real data application, we demonstrate the competitiveness of the proposal empirically. At the same time, the success of the proposed single-scale MOSUM procedure hinges on the availability of the bandwidth $\gamma$ that fulfils a set of assumptions, which may not exist in the presence of multiscale change points (Cho and Kirch,, 2024). One natural avenue for an extension is to apply the MOSUM procedure with a range of bandwidths, which we leave for future research.
Appendix A Proofs
Throughout, we write $C_{NT}=\sqrt{\min(N,T)}$, and denote by $c_{i}\in(0,\infty),\,i\geq 0$, some fixed constants, and by $\epsilon\in(0,1)$ a small constant which may vary from one occasion to another.
A.1 Preliminary lemmas
The following quantities are used extensively in Bai, (2003) and also in our proofs:
$\gamma_{s,t}=\mathsf{E}\left(\frac{1}{N}\sum_{i=1}^{N}e_{i,t}e_{i,s}\right),$ $\qquad\zeta_{s,t}=\frac{1}{N}\sum_{i=1}^{N}e_{i,t}e_{i,s}-\gamma_{s,t},$
$\eta_{s,t}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{g}_{s}^{\top}\bm{\lambda}_{i}e_{i,t},$ $\qquad\xi_{s,t}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{g}_{t}^{\top}\bm{\lambda}_{i}e_{i,s}.$
Then, it holds that
$\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}$ $=\bm{\Phi}_{NT}^{-1}\left(\frac{1}{T}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\gamma_{s,t}+\frac{1}{T}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\zeta_{s,t}+\frac{1}{T}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\eta_{s,t}+\frac{1}{T}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\xi_{s,t}\right),$ (A.1)
where $\text{where \ }\mathbf{H}$ $=\frac{1}{NT}(\bm{\Lambda}^{\top}\bm{\Lambda})(\mathbf{G}^{\top}\widehat{\mathbf{G}})\bm{\Phi}_{NT}^{-1},$ (A.2)
with $\bm{\Phi}_{NT}$ denoting the $r\times r$-diagonal matrix with the $r$ largest eigenvalues of $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ on its diagonal.
Lemma A.1.
Under Assumptions 1 (with $\rho=1$) and 5, we have
$\mathsf{E}\left(\left\|\frac{1}{T}\mathbf{G}^{\top}\mathbf{G}-\bm{\Sigma}_{G}\right\|^{2}\right)\leq c_{0}T^{-1}.$
Proof.
Firstly, we show that for any $j=0,\ldots,R$,
$\mathsf{E}\left(\left\|\frac{1}{k_{j+1}-k_{j}}\sum_{t=k_{j}+1}^{k_{j+1}}\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\bm{\Sigma}_{F}\right\|^{2}\right)\leq c_{0}T^{-1}.$ (A.3)
We begin by showing that Assumption 1 (i) entails that $\{\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\bm{\Sigma}_{F}\}$ is an $\mathcal{L}_{\phi}$-decomposable Bernoulli shift with some $\phi>2$. Without loss of generality, let $r=d=1$ for simplicity. Noting that $f_{t}^{2}=h^{2}(\eta_{t},\eta_{t-1},\ldots)$, consider the construction
$\widetilde{f}_{t,\ell}^{2}=h^{2}(\eta_{t},\ldots,\eta_{t-\ell},\eta_{t-\ell-1}^{\prime},\eta_{t-\ell-2}^{\prime},\ldots),$
where $\{\eta_{t}^{\prime}\}_{t\in\mathbb{Z}}$ is a sequence of i.i.d copies of $\eta_{0}$ independent of $\{\eta_{t}\}_{t\in\mathbb{Z}}$, such that $\eta_{t}\overset{\mathcal{D}}{=}\eta^{\prime}_{t}$. Then, $f_{t}^{2}-\widetilde{f}_{t,\ell}^{2}=(f_{t}+\widetilde{f}_{t,\ell})(f_{t}-\widetilde{f}_{t,\ell})$ whence, using the Cauchy-Schwartz inequality and Minkowski’s inequality,
$\left|f_{t}^{2}-\widetilde{f}_{t,\ell}^{2}\right|_{\phi}\leq\left|f_{t}+\widetilde{f}_{t,\ell}\right|_{2\phi}\left|f_{t}-\widetilde{f}_{t,\ell}\right|_{2\phi}\leq 2\left|f_{t}\right|_{2\phi}\left|f_{t}-\widetilde{f}_{t,\ell}\right|_{2\phi}.$
By Assumption 1 (i), we have $|f_{t}|_{2\phi}<\infty$ provided that $\phi\leq 4$, and also that $|f_{t}-\widetilde{f}_{t,\ell}|_{2\phi}\leq c_{0}\ell^{-a}$ with some $a>2$. Hence, $\{f_{t}^{2}-\mathsf{E}(f_{t}^{2})\}$ is an $\mathcal{L}_{\phi}$-decomposable Bernoulli shift with some $a>2$ such that
$\left|f_{t}^{2}-\widetilde{f}_{t,\ell}^{2}\right|_{\phi}\leq c_{1}\ell^{-a}.$
Then, (A.3) follows from Lemma S2.1 of Aue et al., (2014). This, combined with Assumption 5 and the following observations,
$\left\|\frac{1}{T}\sum_{t=1}^{T}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\bm{\Sigma}_{G}\right\|\leq\sum_{j=0}^{R}(\tau_{j+1}-\tau_{j})\left\|\mathbf{A}_{j}\left(\frac{1}{k_{j+1}-k_{j}}\sum_{t=k_{j}+1}^{k_{j+1}}\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\bm{\Sigma}_{F}\right)\mathbf{A}_{j}^{\top}\right\|$
$\leq\sum_{j=0}^{R}(\tau_{j+1}-\tau_{j})\|\mathbf{A}_{j}\|^{2}\left\|\frac{1}{k_{j+1}-k_{j}}\sum_{t=k_{j}+1}^{k_{j+1}}\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\bm{\Sigma}_{F}\right\|,$
concludes the proof. ∎
Lemma A.2.
Let Assumptions 15 hold, and denote by $\bm{\Phi}\in\mathbb{R}^{r\times r}$ the diagonal matrix containing the eigenvalues of $\bm{\Sigma}_{\Lambda}^{1/2}\bm{\Sigma}_{G}\bm{\Sigma}_{\Lambda}^{1/2}$ on its diagonal. Then, we have
$\mathsf{E}\left(\|\bm{\Phi}_{NT}-\bm{\Phi}\|^{4\rho+\epsilon}\right)\leq c_{0}C_{NT}^{-4\rho-\epsilon},$
where $\rho$ is as in Assumptions 1, 3 and 4.
Proof.
The proof follows from standard arguments, which we briefly summarise. We note that the leading $r$ eigenvalues of $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ are identical to those of $(NT)^{-1}\mathbf{X}^{\top}\mathbf{X}$, and
$\mathbf{X}^{\top}\mathbf{X}=\bm{\Lambda}\mathbf{G}^{\top}\mathbf{G}\bm{\Lambda}^{\top}+\mathbf{E}^{\top}\mathbf{E}+\bm{\Lambda}\mathbf{G}^{\top}\mathbf{E}+\mathbf{E}^{\top}\mathbf{G}\bm{\Lambda}^{\top}.$
Let $\Lambda_{j}\left(\mathbf{A}\right)$ denote the $j$-th eigenvalue, sorted in descending order, of a matrix $\mathbf{A}$. Then,
$\mathsf{E}\left(\|\bm{\Phi}_{NT}-\bm{\Phi}\|^{4\rho+\epsilon}\right)$ $\leq\mathsf{E}\left[\left(\sum_{j=1}^{r}\left|\Lambda_{j}\left(\frac{1}{NT}\mathbf{X}^{\top}\mathbf{X}\right)-\Lambda_{j}(\bm{\Phi})\right|^{2}\right)^{2\rho+\epsilon/2}\right]$
$\leq r^{2\rho+\epsilon/2}\mathsf{E}\left(\sum_{j=1}^{r}\left|\Lambda_{j}\left(\frac{1}{NT}\mathbf{X}^{\top}\mathbf{X}\right)-\Lambda_{j}(\bm{\Phi})\right|^{4\rho+\epsilon}\right).$
By Weyl’s inequality, we have
$\left|\Lambda_{j}\left(\frac{1}{NT}\mathbf{X}^{\top}\mathbf{X}\right)-\Lambda_{j}\left(\frac{1}{NT}\bm{\Lambda}\mathbf{G}^{\top}\mathbf{G}\bm{\Lambda}^{\top}\right)\right|^{4\rho+\epsilon}$
$=$ $\left|\Lambda_{j}\left(\frac{1}{NT}\mathbf{X}^{\top}\mathbf{X}\right)-\Lambda_{j}\left(\frac{1}{NT}\mathbf{G}^{\top}\mathbf{G}\bm{\Lambda}^{\top}\bm{\Lambda}\right)\right|^{4\rho+\epsilon}$
$\leq$ $2\left(\Lambda_{\max}\left(\frac{1}{NT}\mathbf{E}^{\top}\mathbf{E}\right)\right)^{4\rho+\epsilon}+2\left(\Lambda_{\max}\left(\frac{1}{NT}\left(\bm{\Lambda}\mathbf{G}^{\top}\mathbf{E}+\mathbf{E}^{\top}\mathbf{G}\bm{\Lambda}^{\top}\right)\right)\right)^{4\rho+\epsilon}.$
Further
$\Lambda_{\max}\left(\frac{1}{NT}\mathbf{E}^{\top}\mathbf{E}\right)\leq\Lambda_{\max}\left(\frac{1}{NT}\mathsf{E}\left(\mathbf{E}^{\top}\mathbf{E}\right)\right)+\left\|\frac{1}{NT}\left(\mathbf{E}^{\top}\mathbf{E}-\mathsf{E}\left(\mathbf{E}^{\top}\mathbf{E}\right)\right)\right\|.$
Using Assumption 3 (v),
$\Lambda_{\max}\left(\frac{1}{NT}\mathsf{E}\left(\mathbf{E}^{\top}\mathbf{E}\right)\right)\leq\frac{1}{NT}\max_{1\leq i\leq N}\sum_{j=1}^{N}\sum_{s=1}^{T}\left|\mathsf{E}\left(e_{i,t}e_{j,t}\right)\right|\leq c_{0}N^{-1}.$
Similarly, on account of Assumption 3 (vi),
$\mathsf{E}\left(\left\|\frac{1}{NT}\left(\mathbf{E}^{\top}\mathbf{E}-\mathsf{E}\left(\mathbf{E}^{\top}\mathbf{E}\right)\right)\right\|_{F}^{4\rho+\epsilon}\right)$
$=$ $\mathsf{E}\left\{\left[\frac{1}{N^{2}}\sum_{i,j=1}^{N}\left|\frac{1}{T}\sum_{t=1}^{T}\left(e_{i,t}e_{j,t}-\mathsf{E}(e_{i,t}e_{j,t})\right)\right|^{2}\right]^{2\rho+\epsilon/2}\right\}$
$\leq$ $\mathsf{E}\left[\frac{1}{N^{2}}\sum_{i,j=1}^{N}\left|\frac{1}{T}\sum_{t=1}^{T}\left(e_{i,t}e_{j,t}-\mathsf{E}(e_{i,t}e_{j,t})\right)\right|^{4\rho+\epsilon}\right]\leq c_{0}T^{-2\rho-\epsilon/2},$
and therefore
$\left(\Lambda_{\max}\left(\frac{1}{NT}\mathbf{E}^{\top}\mathbf{E}\right)\right)^{4\rho+\epsilon}\leq c_{0}C_{N,T}^{-4\rho-\epsilon}.$
Besides, by Assumption 4 (iii),
$\mathsf{E}\left(\left\|\frac{1}{NT}\bm{\Lambda}\mathbf{G}^{\top}\mathbf{E}\right\|_{F}^{4\rho+\epsilon}\right)$
$=$ $\mathsf{E}\left\{\left[\frac{1}{N^{2}}\sum_{i,j=1}^{N}\left(\frac{1}{T}\sum_{t=1}^{T}\bm{\lambda}_{i}^{\top}\mathbf{g}_{t}e_{j,t}\right)^{2}\right]^{2\rho+\epsilon/2}\right\}$
$\leq$ $\frac{1}{N^{2}}\sum_{i,j=1}^{N}\mathsf{E}\left[\left(\frac{1}{T}\sum_{t=1}^{T}\bm{\lambda}_{i}^{\top}\mathbf{g}_{t}e_{j,t}\right)^{4\rho+\epsilon}\right]\leq c_{0}T^{-2\rho-\epsilon/2}$
and $\mathsf{E}(\|(NT)^{-1}\mathbf{E}^{\top}\mathbf{G}\bm{\Lambda}^{\top}\|_{F}^{4\rho+\epsilon})$ is similarly bounded. Finally, denoting by $\mathbf{B}=\bm{\Sigma}_{\Lambda}^{1/2}\bm{\Sigma}_{G}\bm{\Sigma}_{\Lambda}^{1/2}$ and $\mathbf{B}_{NT}=(NT)^{-1}(\bm{\Lambda}^{\top}\bm{\Lambda})^{1/2}(\mathbf{G}^{\top}\mathbf{G})(\bm{\Lambda}^{\top}\bm{\Lambda})^{1/2}$, we have $\|\mathbf{B}_{NT}-\mathbf{B}\|=O_{P}(C_{NT}^{-1})$ by Assumption 2 (ii) and Lemma A.1. The desired result now follows from putting all the bounds together. ∎
Bai, (2003) proves that $T^{-1}\sum_{t=1}^{T}\|\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\|^{2}$ and $T^{-1}\|\sum_{t=1}^{T}\mathbf{g}_{t}(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t})^{\top}\|$ are bounded in probability. We report two results of independent interest by deriving the upper bounds on the two terms in $\mathcal{L}_{\delta}$-norm for some $\delta\geq 1$.
Lemma A.3.
Suppose that Assumptions 1 and 3 hold with $\rho=1$. Then it follows that
$\mathsf{E}\left(\left\|\sum_{t=1}^{T}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|^{\delta}\right)\leq c_{0}\left(TC_{NT}^{-2}\right)^{\delta}$
for all $1\leq\delta\leq 2+\epsilon$, with $\mathbf{H}$ defined in (A.2).
Proof.
To simplify the notation, we set $r=d=1$ and omit the matrix $\mathbf{H}$. By convexity,
$\mathsf{E}\left(\left|\frac{1}{T}\sum_{t=1}^{T}\left(\widehat{g}_{t}-g_{t}\right)^{2}\right|^{\delta}\right)\leq\frac{1}{T}\sum_{t=1}^{T}\mathsf{E}\left(\left|\widehat{g}_{t}-g_{t}\right|^{2\delta}\right).$
Using (A.1),
$\frac{1}{T}\sum_{t=1}^{T}\mathsf{E}\left(\left|\widehat{g}_{t}-g_{t}\right|^{2\delta}\right)$ $\leq T^{-2\delta-1}\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t}\right|^{2\delta}\right)+T^{-2\delta-1}\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t}\right|^{2\delta}\right)$
$+T^{-2\delta-1}\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t}\right|^{2\delta}\right)+T^{-2\delta-1}\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\widehat{g}_{s}\xi_{s,t}\right|^{2\delta}\right)$
$=:T_{1}+T_{2}+T_{3}+T_{4}.$ (A.4)
We now study each of these terms. By construction, $\sum_{s=1}^{T}\widehat{g}_{s}^{2}=T$. Also by Assumption 3 (ii) and Lemma 1 (i) in Bai and Ng, (2002), it follows that $\sum_{s=1}^{T}\gamma_{s,t}^{2}\leq c_{0}$. Therefore,
$T_{1}\leq\mathsf{E}\left(\sum_{t=1}^{T}\left|\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right|^{\delta}\left|\sum_{s=1}^{T}\gamma_{s,t}^{2}\right|^{\delta}\right)\leq c_{0}T^{-\delta}.$ (A.5)
Next, from Assumption 3 (iii),
$T^{2\delta+1}\cdot T_{2}$ $\leq\sum_{t=1}^{T}\mathsf{E}\left(\left|\left(\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right)^{1/2}\left(\sum_{s=1}^{T}\zeta_{s,t}^{2}\right)^{1/2}\right|^{2\delta}\right)$
$\leq T^{\delta}\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\zeta_{s,t}^{2}\right|^{\delta}\right)\leq T^{2\delta-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathsf{E}\left(\left|\zeta_{s,t}\right|^{2\delta}\right)\leq T^{2\delta+1}N^{-\delta},$
Therefore it follows that
$T_{2}\leq c_{0}N^{-\delta}.$ (A.6)
Similarly, by the Cauchy-Schwartz inequality,
$T^{2\delta+1}\cdot T_{3}$ $=\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\widehat{g}_{s}g_{s}\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{2\delta}\right)$
$\leq\sum_{t=1}^{T}\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{2\delta}\left|\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right|^{\delta}\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta}\right)$
$=T^{\delta}\sum_{t=1}^{T}\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{2\delta}\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta}\right)$
$\leq T^{\delta}\sum_{t=1}^{T}\left[\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{4\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{2\delta}\right)\right]^{1/2}.$
By Assumption 3 (iv),
$\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{4\delta}\right)\leq c_{0}N^{-2\delta}.$
Also, from Assumption 1 (i),
$\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{2\delta}\right)\leq T^{2\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|g_{s}\right|^{4\delta}\right)\leq c_{0}T^{2\delta}.$
Putting all together, we have
$T_{3}\leq c_{0}N^{-\delta}.$ (A.7)
Finally, we consider
$T^{2\delta+1}\cdot T_{4}$ $\leq\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right|^{\delta}\left|\sum_{s=1}^{T}\xi_{s,t}^{2}\right|^{\delta}\right)$
$=T^{\delta}\sum_{t=1}^{T}\mathsf{E}\left(\left|\sum_{s=1}^{T}\xi_{s,t}^{2}\right|^{\delta}\right)\leq T^{2\delta-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathsf{E}\left(\left|\xi_{s,t}\right|^{2\delta}\right).$
By Assumptions 3 (iv) and 1 (i),
$\mathsf{E}\left(\left|\xi_{s,t}\right|^{2\delta}\right)=\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}g_{t}\lambda_{i}e_{i,s}\right|^{2\delta}\right)\leq\left[\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{4\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|g_{t}\right|^{4\delta}\right)\right]^{1/2}\leq c_{0}N^{-\delta},$
Therefore it follows that
$T_{4}\leq c_{0}N^{-\delta}.$ (A.8)
Putting together (A.5)–(A.8) into (A.4), the desired result follows. ∎
Lemma A.4.
Suppose that Assumptions 15 hold. Then it holds that
$\mathsf{E}\left(\left\|\sum_{t=1}^{T}\mathbf{g}_{t}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|^{\delta}\right)\leq c_{0}\left(TC_{NT}^{-2}\right)^{\delta},$
for all $1\leq\delta\leq\rho+\epsilon$, with $\rho$ is as in Assumptions 1, 3 and 4.
Proof.
Throughout, we frequently use that for $1\leq\delta\leq 4+\epsilon$,
$\mathsf{E}\left(\left\|\sum_{s=1}^{T}\mathbf{g}_{s}\mathbf{g}_{s}^{\top}\right\|^{\delta}\right)\leq T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\|\mathbf{g}_{s}\|^{2\delta}\right)\leq c_{0}T^{\delta},$ (A.9)
from Assumption 1 (i). Thanks to (A.1), we have
$\mathsf{E}\left(\left\|\sum_{t=1}^{T}\mathbf{g}_{t}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|^{\delta}\right)$ $\leq\mathsf{E}\left(\left\|\frac{1}{T}\sum_{s,t=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\gamma_{s,t}\right\|^{\delta}\right)+\mathsf{E}\left(\left\|\frac{1}{T}\sum_{s,t=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\zeta_{s,t}\right\|^{\delta}\right)$
$+\mathsf{E}\left(\left\|\frac{1}{T}\sum_{s,t=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\eta_{s,t}\right\|^{\delta}\right)+\mathsf{E}\left(\left\|\frac{1}{T}\sum_{s,t=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)$
$=:T_{1}+T_{2}+T_{3}+T_{4}.$ (A.10)
We begin with $T_{4}$ which is bounded as
$T_{4}\leq T^{-\delta}\mathsf{E}\left(\left\|\sum_{s,t=1}^{T}(\widehat{\mathbf{g}}_{s}-\mathbf{H}^{\top}\mathbf{g}_{s})\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)+T^{-\delta}\mathsf{E}\left(\left\|\mathbf{H}^{\top}\sum_{s,t=1}^{T}\mathbf{g}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)=:T_{4,1}+T_{4,2}.$
By Assumptions 1, 2 and 5 and Lemma A.2, we have
$\mathsf{E}\left(\left\|\mathbf{H}\right\|^{p\delta}\right)$ $\leq\left\|\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right\|^{p\delta}\mathsf{E}\left(\left\|\frac{1}{\sqrt{T}}\widehat{\mathbf{G}}\right\|^{p\delta}\left\|\frac{1}{\sqrt{T}}\mathbf{G}\right\|^{p\delta}\|\bm{\Phi}_{NT}^{-1}\|^{p\delta}\right)$
$\leq c_{0}\left\{\mathsf{E}\left[\left(\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{g}_{t}\|^{2}\right)^{p\delta}\right]\;\mathsf{E}\left(\|\bm{\Phi}_{NT}^{-1}\|^{2p\delta}\right)\right\}^{1/2}$
$\leq c_{1}\left\{\frac{1}{T}\sum_{t=1}^{T}\mathsf{E}\left(\left\|\mathbf{g}_{t}\right\|^{2p\delta}\right)\right\}^{1/2}\leq c_{2}$ (A.11)
for $1\leq p\leq 2$. From this, we obtain
$T^{\delta}\cdot T_{4,2}$ $\leq\mathsf{E}\left(\|\mathbf{H}^{\top}\|^{\delta}\left\|\sum_{s,t=1}^{\top}\mathbf{g}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)$
$\leq\left[\mathsf{E}\left(\|\mathbf{H}^{\top}\|^{2\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left\|\sum_{s,t=1}^{\top}\mathbf{g}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{2\delta}\right)\right]^{1/2}.$
WLOG, we may set $r=d=1$ for notational simplicity. Then by Hölder’s inequality,
$\mathsf{E}\left(\left|\sum_{s,t=1}^{T}g_{s}g_{t}\xi_{s,t}\right|^{2\delta}\right)$ $=\mathsf{E}\left[\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{2\delta}\left|\frac{1}{N}\sum_{s=1}^{T}\sum_{i=1}^{N}\lambda_{i}g_{s}e_{i,s}\right|^{2\delta}\right]$
$\leq\left(\mathsf{E}\left[\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{4\delta}\right]\right)^{1/2}\left(\mathsf{E}\left[\left|\frac{1}{N}\sum_{s=1}^{T}\sum_{i=1}^{N}\lambda_{i}g_{s}e_{i,s}\right|^{4\delta}\right]\right)^{1/2},$
where by (A.9),
$\left(\mathsf{E}\left[\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{4\delta}\right]\right)^{1/2}\leq\left[T^{4\delta-1}\sum_{t=1}^{T}\mathsf{E}(|g_{t}|^{8\delta})\right]^{1/2}\leq c_{0}T^{2\delta},$
while by Assumption 4 (ii),
$\left(\mathsf{E}\left[\left|\frac{1}{N}\sum_{s=1}^{T}\sum_{i=1}^{N}\lambda_{i}g_{s}e_{i,s}\right|^{4\delta}\right]\right)^{1/2}\leq c_{0}(T^{1/2}N^{-1/2})^{2\delta}.$
Altogether, this yields
$T_{4,2}\leq c_{0}T^{\delta/2}N^{-\delta/2}.$ (A.12)
Next, we note that by Hölder’s inequality with some $1<p\leq 4$,
$T_{4,1}$ $=T^{-\delta}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}^{2}\cdot\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{\delta}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}^{2}\right|^{p\delta}\right)\right]^{\frac{1}{p}}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{\frac{p\delta}{p-1}}\right)\right]^{\frac{p-1}{p}}$
$\leq c_{0}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{\frac{p\delta}{p-1}}\right)\right]^{\frac{p-1}{p}},$
where the last passage follows from (A.9). Again applying Hölder’s inequality,
$\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\left(\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right)\right|^{\frac{p\delta}{p-1}}\right)$
$\leq\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\frac{p\delta}{2(p-1)}}\left|\sum_{s=1}^{T}\left(\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right)^{2}\right|^{\frac{p\delta}{2(p-1)}}\right)$
$\leq\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\frac{pq\delta}{2(p-1)}}\right)\right]^{\frac{1}{q}}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right)^{2}\right|^{\frac{pq\delta}{2(p-1)(q-1)}}\right)\right]^{\frac{q-1}{q}}.$
Setting $p=4$ and $q=3/2$, we have $pq\delta/(p-1)=2\delta$ such that by Lemma A.3,
$\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\frac{pq\delta}{2(p-1)}}\right)\leq c_{0}\left(TC_{NT}^{-2}\right)^{\delta}.$
Besides, since the choices of $p$ and $q$ lead to
$\frac{pq\delta}{2(p-1)(q-1)}=2\delta,$
applying Assumption 3 (iv),
$\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right)^{2}\right|^{2\delta}\right)$ $\leq T^{2\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{4\delta}\right)$
$\leq c_{0}T^{2\delta}N^{-2\delta}.$
From the above arguments and (A.9), it follows that
$T_{4,1}\leq c_{0}T^{\delta}N^{-\delta/2}C_{NT}^{-\delta}.$ (A.13)
Collecting the bounds on $T_{4,1}$ and $T_{4,2}$, we have
$T_{4}\leq c_{0}T^{\delta}N^{-\delta/2}C_{NT}^{-\delta}.$ (A.14)
For the rest of the terms, we may proceed analogously. From (A.11), for simplicity, we check the steps by setting $r=d=1$ and omitting $\mathbf{H}$. Note that
$T_{1}\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s,t=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)g_{t}\gamma_{s,t}\right|^{\delta}\right)+T^{-\delta}\mathsf{E}\left(\left|\sum_{s,t=1}^{T}g_{s}g_{t}\gamma_{s,t}\right|^{\delta}\right)=:T_{1,1}+T_{1,2}.$
Using the Cauchy-Schwartz inequality twice,
$T_{1,1}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta/2}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right|^{2\delta}\right)\right]^{1/2},$
having used Lemma A.3 in the last passage. Let $n_{0}=\lceil 4\delta\rceil$ and recall that, by Assumption 1 (i), $\mathsf{E}(\left|g_{t}\right|^{n_{0}})<\infty$. Using the $\mathcal{L}_{p}$-norm inequality, we have
$\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right|^{2\delta}\right)$ $\leq\left[\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right|^{n_{0}}\right)\right]^{2\delta/n_{0}},\text{ \ and}$
$\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right|^{n_{0}}\right)$ $=\mathsf{E}\left(\left|\sum_{t_{1},\ldots,t_{n_{0}}=1}^{T}g_{t_{1}}\cdot\ldots\cdot g_{t_{n_{0}}}\cdot\gamma_{s,t_{1}}\cdot\ldots\cdot\gamma_{s,t_{n_{0}}}\right|\right)$
$\leq\sum_{t_{1},\ldots,t_{n_{0}}=1}^{T}\mathsf{E}\left(\prod\limits_{i=1}^{n_{0}}\left|g_{t_{i}}\right|\right)\prod\limits_{i=1}^{n_{0}}\left|\gamma_{s,t_{i}}\right|$
$\leq\max_{1\leq t\leq T}\left|g_{t_{i}}\right|_{n_{0}}\cdot\left(\sum_{s=1}^{T}\left|\gamma_{s,t}\right|\right)^{n_{0}}\leq c_{0},$
where the penultimate passage follows from Hölder’s inequality, and the last passage from Assumption 3 (ii). The above entails that $\mathsf{E}(|\sum_{t=1}^{T}g_{t}\gamma_{s,t}|^{2\delta})\leq c_{0}$, and therefore $T_{1,1}\leq c_{0}C_{NT}^{-\delta}$. Similarly, we have from (A.9),
$T_{1,2}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta/2}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta}\right)\right]^{1/2}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\gamma_{s,t}\right|^{2\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta}\cdot T^{\delta/2}\cdot T^{\delta/2}=c_{0},$
so that we finally have
$T_{1}\leq c_{0}.$ (A.15)
We now turn to studying
$T_{2}\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\left(\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right)\right|^{\delta}\right)+T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}\left(\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right)\right|^{\delta}\right)=:T_{2,1}+T_{2,2}.$
Using the Cauchy-Schwartz inequality and Lemma A.3,
$T_{2,1}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta/2}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right|^{2\delta}\right)\right]^{1/2};$
thence, Assumption 4 (i) immediately entails that
$T_{2,1}\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left(T^{\delta}\left(N^{-1}T\right)^{\delta}\right)^{1/2}=c_{0}T^{\delta/2}N^{-\delta/2}C_{NT}^{-\delta}.$
By the same token, with (A.9),
$T_{2,2}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta/2}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\zeta_{s,t}\right|^{2\delta}\right)\right]^{1/2}\leq c_{0}T^{\delta/2}N^{-\delta/2}.$
Putting together the bounds on $T_{2,1}$ and $T_{2,2}$, we have
$T_{2}\leq c_{0}T^{\delta/2}N^{-\delta/2}.$ (A.16)
Finally consider
$T_{3}\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\left(\sum_{t=1}^{T}g_{t}\eta_{s,t}\right)\right|^{\delta}\right)+T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}\left(\sum_{t=1}^{T}g_{t}\eta_{s,t}\right)\right|^{\delta}\right)=:T_{3,1}+T_{3,2}.$
It holds that
$T_{3,1}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta/2}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\eta_{s,t}\right|^{2\delta}\right)\right]^{1/2}.$
By definition of $\eta_{s,t}$, making use of (A.9),
$T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\eta_{s,t}\right|^{2\delta}\right)$ $=T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}g_{s}\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{2\delta}\right)$
$\leq T^{\delta-1}\sum_{s=1}^{T}\left[\mathsf{E}\left(\left|g_{s}\right|^{4\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}g_{t}\lambda_{i}e_{i,t}\right|^{4\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{2\delta}N^{-\delta}$
by Assumptions 1 (i) and 4 (ii). Therefore,
$T_{3,1}\leq c_{0}T^{\delta/2}N^{-\delta/2}C_{NT}^{-\delta}.$
Along the same lines,
$T_{3,2}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta/2}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{T}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}\eta_{s,t}\right|^{2\delta}\right)\right]^{1/2}\leq c_{0}T^{\delta/2}N^{-\delta/2},$
which entails that
$T_{3}\leq c_{0}T^{\delta/2}N^{-\delta/2}.$ (A.17)
The desired result now follows from plugging (A.14), (A.15), (A.16) and (A.17) into (A.10). ∎
Lemma A.5.
Suppose that Assumptions 15 hold with $\rho=1$ in Assumptions 1, 3 and 4. Then it follows that
$\left\|\left(\frac{\widehat{\mathbf{G}}^{\top}\mathbf{G}}{T}\right)\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)\left(\frac{\mathbf{G}^{\top}\widehat{\mathbf{G}}}{T}\right)-\bm{\Phi}\right\|=O_{P}\left(\frac{1}{C_{NT}}\right).$
Proof.
By construction, we have
$\widehat{\mathbf{G}}\bm{\Phi}_{NT}$ $=\left(\frac{\mathbf{X}\mathbf{X}^{\top}}{NT}\right)\widehat{\mathbf{G}},\text{ \ hence}$ (A.18)
$\bm{\Phi}_{NT}$ $=\frac{1}{T}\widehat{\mathbf{G}}^{\top}\left(\frac{\mathbf{X}\mathbf{X}^{\top}}{NT}\right)\widehat{\mathbf{G}}=\frac{\widehat{\mathbf{G}}^{\top}\mathbf{G}}{T}\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)\frac{\mathbf{G}^{\top}\widehat{\mathbf{G}}}{T}+\frac{1}{T}\widehat{\mathbf{G}}^{\top}\mathbf{R}_{NT}\widehat{\mathbf{G}},\text{ \ where}$
$\mathbf{R}_{NT}$ $=\frac{1}{NT}\left(\mathbf{G}\bm{\Lambda}^{\top}\mathbf{E}^{\top}+\mathbf{E}\bm{\Lambda}\mathbf{G}^{\top}+\mathbf{E}\mathbf{E}^{\top}\right).$
By (A.1) and Lemmas A.3 and A.4,
$\left\|\frac{1}{T}\widehat{\mathbf{G}}^{\top}\mathbf{R}_{NT}\widehat{\mathbf{G}}\right\|$ $\leq\left(\left\|\frac{1}{T}(\widehat{\mathbf{G}}-\mathbf{G}\mathbf{H})^{\top}(\widehat{\mathbf{G}}-\mathbf{G}\mathbf{H})\right\|+\left\|\mathbf{H}\right\|\;\left\|\frac{1}{T}\mathbf{G}^{\top}(\widehat{\mathbf{G}}-\mathbf{G}\mathbf{H})\right\|\right)\left\|\bm{\Phi}_{NT}^{-1}\right\|$
$=O_{P}\left(\frac{1}{C_{NT}^{2}}\right),$
where we also use that $\|\bm{\Phi}_{NT}^{-1}\|=O_{P}(1)$ from Lemma A.2, and $\|\mathbf{H}\|=O_{P}(1)$ from (A.11). Then, the conclusion follows from Lemma A.2. ∎
Lemma A.6.
Suppose that Assumptions 15 hold with $\rho=1$ in Assumptions 1, 3 and 4. For $\mathbf{H}_{0}$ is defined in (9), we have $\|\mathbf{H}-\mathbf{H}_{0}\|=O_{P}(C_{NT}^{-1})$.
Proof.
From (A.18), $\widehat{\mathbf{G}}$ satisfies
$\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)^{1/2}\frac{1}{T}\mathbf{G}^{\top}\left(\frac{\mathbf{X}\mathbf{X}^{\top}}{NT}\right)\widehat{\mathbf{G}}$ $=\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)^{1/2}\left(\frac{\mathbf{G}^{\top}\widehat{\mathbf{G}}}{T}\right)\bm{\Phi}_{NT}.$
Substituting $\mathbf{X}=\mathbf{G}\bm{\Lambda}^{\top}+\mathbf{E}$ into the above equations, we have
$\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)^{1/2}\left(\frac{\mathbf{G}^{\top}\mathbf{G}}{T}\right)\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)\left(\frac{\mathbf{G}^{\top}\widehat{\mathbf{G}}}{T}\right)+c_{NT}$ $=\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)^{1/2}\left(\frac{\mathbf{G}^{\top}\widehat{\mathbf{G}}}{T}\right)\bm{\Phi}_{NT}$
where, recalling the definition of $\mathbf{R}_{NT}$ in the proof of Lemma A.5, we have
$c_{NT}$ $=\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)^{1/2}\frac{1}{T}\mathbf{G}^{\top}\mathbf{R}_{NT}\widehat{\mathbf{G}}\text{ \ such that \ }$
$\|c_{NT}\|$ $\leq\left\|\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right\|^{1/2}\left\|\frac{1}{T}\mathbf{G}^{\top}(\widehat{\mathbf{G}}-\mathbf{G}\mathbf{H})\right\|=O_{P}\left(\frac{1}{C_{NT}^{2}}\right).$
by Assumption 2 (ii) and Lemma A.4. Recall the definitions of $\mathbf{B}$ and $\mathbf{B}_{NT}$ in the proof of Lemma A.2, and let us define
$\mathbf{C}_{NT}=\left(\frac{\bm{\Lambda}^{\top}\bm{\Lambda}}{N}\right)^{1/2}\left(\frac{\mathbf{G}^{\top}\widehat{\mathbf{G}}}{T}\right).$
Then, $\|\mathbf{B}_{NT}-\mathbf{B}\|=O_{P}(C_{NT}^{-1})$ by Assumption 2 (ii) and Lemma A.1. Also, $\mathbf{C}_{NT}$ is $O_{P}(1)$ and (asymptotically) invertible from Lemma A.5. Denote by $\mathbf{W}$ the $r\times r$-matrix containing the (normalised) eigenvectors of $\mathbf{B}$ corresponding to the eigenvalues on the diagonal of $\bm{\Phi}$. Then, the remainder of the proof proceeds analogously as that of Lemma 6 of Han and Inoue, (2015) which shows that $\mathbf{H}_{0}=\text{plim}_{\min(N,T)\rightarrow\infty}\mathbf{H}=\bm{\Sigma}_{\Lambda}^{1/2}\mathbf{W}\bm{\Phi}^{-1/2}$ and $\|\mathbf{H}-\mathbf{H}_{0}\|=O_{P}(C_{NT}^{-1})$. ∎
The next two lemmas contain two maximal inequalities which are required to bound the difference between the partial sums of $\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}$ and those of $\mathbf{H}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}$.
Lemma A.7.
Suppose that the assumptions of Lemma A.3 hold. Then, we have
$\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|^{\delta}\right)\leq c_{0}T^{\delta}C_{NT}^{-2\delta},$
for all $1\leq\delta\leq 2+\epsilon$.
Proof.
The proof of follows immediately upon noting that
$\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|^{\delta}\right)\leq\mathsf{E}\left(\left\|\sum_{t=1}^{T}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|^{\delta}\right)$
and using Lemma A.3. ∎
Lemma A.8.
Suppose that the assumptions of Lemma A.4 hold with $\rho=1$ in Assumptions 1, 3 and 4. Then,
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\gamma_{s,t}\right\|^{\delta}\right)$ $\leq c_{0},$ (A.19)
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\zeta_{s,t}\right\|^{\delta}\right)$ $\leq c_{0}T^{{\delta/2}}N^{-\delta/2},$ (A.20)
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\eta_{s,t}\right\|^{\delta}\right)$ $\leq c_{0}T^{\delta_{2}/2}N^{-\delta/2},$ (A.21)
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)$ $\leq c_{0}T^{\delta_{3}/2}N^{-\delta/2}+c_{1}T^{\delta}N^{-\delta/2}{C_{NT}^{-\delta}},$ (A.22)
for $1\leq\delta\leq 1+\epsilon$.
Proof.
The proof is based on very similar passages as the proof of Lemma A.4, which we omit when possible. As before, we start with (A.22). Note that
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{\mathbf{g}}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)\leq T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}(\widehat{\mathbf{g}}_{s}-\mathbf{H}^{\top}\mathbf{g}_{s})\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)$
$+T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\mathbf{H}^{\top}\sum_{t=1}^{k}\sum_{s=1}^{T}\mathbf{g}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{\delta}\right)=:T_{4,1}+T_{4,2}.$
Repeating the same passages as in the proof of Lemma A.4 leading to (A.12), it is easily seen that
$T_{4,2}$ $\leq T^{-\delta}\left(\mathsf{E}(\|\mathbf{H}^{\top}\|^{2\delta}\right)^{1/2}\left[\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\sum_{s=1}^{T}\mathbf{g}_{s}\mathbf{g}_{t}^{\top}\xi_{s,t}\right\|^{2\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta}\left\{\left[\mathsf{E}\left(\max_{1\leq k\leq T}\left\|\sum_{t=1}^{k}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right\|^{4\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left\|\frac{1}{N}\sum_{s=1}^{T}\sum_{i=1}^{N}\bm{\lambda}_{i}^{\top}\mathbf{g}_{s}e_{i,s}\right\|^{4\delta}\right)\right]^{1/2}\right\}^{1/2}$
$\leq c_{0}T^{-\delta}\left\{\left[\mathsf{E}\left(\left\|\sum_{t=1}^{T}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right\|^{4\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\left\|\frac{1}{N}\sum_{s=1}^{T}\sum_{i=1}^{N}\bm{\lambda}_{i}^{\top}\mathbf{g}_{s}e_{i,s}\right\|^{4\delta}\right)\right]^{1/2}\right\}^{1/2}$
$\leq c_{0}T^{\delta/2}N^{-\delta/2}$
As for $T_{4,1}$, we may set $r=d=1$ and omit $\mathbf{H}$, which gives
$T_{4,1}$ $=T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}^{2}\cdot\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{\delta}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{t=1}^{T}g_{t}^{2}\right|^{p\delta}\right)\right]^{\frac{1}{p}}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,s}\right|^{\frac{p\delta}{p-1}}\right)\right]^{\frac{p-1}{p}}.$
Setting $p=4$ and applying the arguments analogous to those adopted in (A.13), we obtain $T_{4,1}\leq c_{0}T^{\delta}N^{-\delta/2}C_{NT}^{-\delta}$ which completes the proof of (A.22).
For the rest of the proof, we proceed analogously and check the steps for the case of $r=d=1$ for simplicity. For (A.19), we have
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{g}_{s}g_{t}\gamma_{s,t}\right|^{\delta}\right)\leq T^{-\delta}\mathsf{E}\left(\left|\max_{1\leq k\leq T}\sum_{t=1}^{k}\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)g_{t}\gamma_{s,t}\right|^{\delta}\right)$
$+T^{-\delta}\mathsf{E}\left(\left|\max_{1\leq k\leq T}\sum_{t=1}^{k}\sum_{s=1}^{T}g_{s}g_{t}\gamma_{s,t}\right|^{\delta}\right)=:T_{1,1}+T_{1,2}.$
Using the Cauchy-Schwartz inequality twice,
$T_{1,1}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta/2}\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\gamma_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\gamma_{s,t}\right|^{2\delta}\right)\right]^{1/2},$
having used Lemma A.3 in the last passage. Let $n_{0}=\lceil 4\delta\rceil$. Using the $\mathcal{L}_{p}$-norm inequality, we have
$\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\gamma_{s,t}\right|^{n_{0}}\right)$ $=\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t_{1},\ldots,t_{n_{0}}=1}^{k}g_{t_{1}}\cdot\ldots\cdot g_{t_{n_{0}}}\cdot\gamma_{s,t_{1}}\cdot\ldots\cdot\gamma_{s,t_{n_{0}}}\right|\right)$
$\leq\sum_{t_{1},\ldots,t_{n_{0}}=1}^{T}\mathsf{E}\left(\prod\limits_{i=1}^{n_{0}}\left|g_{t_{i}}\right|\right)\prod\limits_{i=1}^{n_{0}}\left|\gamma_{s,t_{i}}\right|\leq c_{0},$
and therefore $T_{1,1}\leq c_{0}C_{NT}^{-\delta}$. Analgously, combined with (A.9),
$T_{1,2}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta/2}\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\gamma_{s,t}\right)^{2}\right|^{\delta/2}\right)\leq c_{0}T^{-\delta}\cdot T^{\delta/2}\cdot T^{\delta/2}=O(1),$
so we have the desired result. Next, let us consider (A.20).
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{g}_{s}g_{t}\zeta_{s,t}\right|^{\delta}\right)\leq T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\left(\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right)\right|^{\delta}\right)$
$+T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}g_{s}\left(\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right)\right|^{\delta}\right)=:T_{2,1}+T_{2,2}.$
Using the Cauchy-Schwartz inequality and Lemma A.3,
$T_{2,1}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta/2}\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right|^{2\delta}\right)\right]^{1/2}.$
Under Assumption 4 (i), it follows that
$\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right|^{2\delta}\right)\leq c_{0}T^{\delta}N^{-\delta}$
by Theorem 3.1 in Móricz et al., (1982), which immediately entails that
$T_{2,1}\leq c_{0}T^{\delta/2}N^{-\delta/2}C_{NT}^{-\delta}.$ (A.23)
By the same token, with (A.9),
$T_{2,2}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta/2}\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq c_{0}T^{-\delta/2}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\zeta_{s,t}\right|^{2\delta}\right)\right]^{1/2}\leq c_{0}T^{\delta/2}N^{-\delta/2}.$
Putting together the bounds on $T_{2,1}$ and $T_{2,2}$, we have the desired result. Finally, as for (A.21),
$T^{-\delta}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\sum_{t=1}^{k}\widehat{g}_{s}g_{t}\eta_{s,t}\right|^{\delta}\right)\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)\max_{1\leq k\leq T}\left(\sum_{t=1}^{k}g_{t}\eta_{s,t}\right)\right|^{\delta}\right)$
$+T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}\max_{1\leq k\leq T}\left(\sum_{t=1}^{k}g_{t}\eta_{s,t}\right)\right|^{\delta}\right)=:T_{3,1}+T_{3,2}.$
It holds that
$T_{3,1}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta/2}\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}\left(\widehat{g}_{s}-g_{s}\right)^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}C_{NT}^{-\delta}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\eta_{s,t}\right|^{2\delta}\right)\right]^{1/2}.$
By definition of $\eta_{s,t}$, making use of (A.9),
$T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\eta_{s,t}\right|^{2\delta}\right)=T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}g_{s}\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t}\right|^{2\delta}\right)$
$\leq T^{\delta-1}\sum_{s=1}^{T}\left[\mathsf{E}\left(\left|g_{s}\right|^{4\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\max_{1\leq k\leq T}\left|\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{k}g_{t}\lambda_{i}e_{i,t}\right|^{4\delta}\right)\right]^{1/2}\leq c_{0}T^{2\delta}N^{-\delta}$
by Assumptions 1 (i) and 4 (ii), and Theorem 3.1 of Móricz et al., (1982). Therefore,
$T_{3,1}\leq c_{0}T^{\delta/2}N^{-\delta/2}C_{NT}^{-\delta}.$
Along the same lines,
$T_{3,2}$ $\leq T^{-\delta}\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta/2}\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta/2}\right)$
$\leq T^{-\delta}\left[\mathsf{E}\left(\left|\sum_{s=1}^{T}g_{s}^{2}\right|^{\delta}\right)\right]^{1/2}\left[\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{s=1}^{T}\left(\sum_{t=1}^{k}g_{t}\eta_{s,t}\right)^{2}\right|^{\delta}\right)\right]^{1/2}$
$\leq c_{0}T^{-\delta/2}\left[T^{\delta-1}\sum_{s=1}^{T}\mathsf{E}\left(\max_{1\leq k\leq T}\left|\sum_{t=1}^{k}g_{t}\eta_{s,t}\right|^{2\delta}\right)\right]^{1/2}\leq c_{0}T^{\delta/2}N^{-\delta/2},$
which completes the proof. ∎
The following three lemmas provide an estimate of the rate of convergence of the partial sums of (appropriately centered) $\mathsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ to their weak limit.
Lemma A.9.
[…]
$\max_{1\leq k\leq\Delta_{j}/2}\frac{1}{k^{\zeta_{1}}}\left\|\sum_{t=k_{j}+1}^{k_{j}+k}\mathsf{Vech}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\right)-\mathbf{D}_{j}^{1/2}W_{1,dT}^{(j)}(k)\right\|=O_{P}(1),$
$\max_{\Delta_{j}/2<k<\Delta_{j}}\frac{1}{(\Delta_{j}-k)^{\zeta_{1}}}\left\|\sum_{t=k_{j}+k+1}^{k_{j+1}}\mathsf{Vech}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\right)-\mathbf{D}_{j}^{1/2}W_{2,dT}^{(j)}\left(\Delta_{j}-k\right)\right\|=O_{P}(1)$
where, with $\mathbf{A}_{j}$ defined in (2) and $\mathbf{D}$ in (6), we have
$\mathbf{D}_{j}=\mathbf{L}_{r}(\mathbf{A}_{j}\otimes\mathbf{A}_{j})\mathbf{K}_{r}\mathbf{D}\mathbf{K}_{r}^{\top}(\mathbf{A}_{j}^{\top}\otimes\mathbf{A}_{j}^{\top})\mathbf{L}_{r}^{\top}.$ (A.24)
Proof.
[…]
$\max_{1\leq k\leq T/2}\frac{1}{k^{\zeta_{1}}}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\mathsf{E}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\top}\right)\right)-\mathbf{D}^{1/2}W_{1,dT}(k)\right\|=O_{P}(1),$
$\max_{T/2<k<T}\frac{1}{(T-k)^{\zeta_{1}}}\left\|\sum_{t=k+1}^{T}\mathsf{Vech}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\mathsf{E}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\top}\right)\right)-\mathbf{D}^{1/2}W_{2,dT}(T-k)\right\|=O_{P}(1).$
[…]
Lemma A.10.
Suppose that the assumptions of Lemmas A.3 and A.4 hold with with $\rho=1$ in Assumptions 1, 3 and 4, as well as Assumption 6. Then there exists some constant $\zeta_{2}\in(0,1/2)$ such that
$\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|$ $=O_{P}(1),$ (A.25)
$\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left\|\sum_{t=1}^{k}\mathbf{g}_{t}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|$ $=O_{P}(1).$ (A.26)
Proof.
We begin with (A.25). Standard arguments entail that
$\mathsf{P}\left(\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|>x\right)$
$\leq$ $\,\mathsf{P}\left(\max_{0\leq\ell\leq\lfloor\log(T)\rfloor}\max_{\exp(\ell)\leq k\leq\exp(\ell+1)}\frac{1}{k^{\zeta_{2}}}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|>x\right)$
$\leq$ $\,\sum_{\ell=0}^{\lfloor\log(T)\rfloor}\mathsf{P}\left(\max_{\exp(\ell)\leq k\leq\exp(\ell+1)}\frac{1}{k^{\zeta_{2}}}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|>x\right)$
$\leq$ $\,\sum_{\ell=0}^{\lfloor\log(T)\rfloor}\mathsf{P}\left(\max_{\exp(\ell)\leq k\leq\exp(\ell+1)}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|>x\exp(\zeta_{2}\ell)\right)$
$\leq$ $\,\frac{1}{x}\sum_{\ell=0}^{\lfloor\log(T)\rfloor}\exp(-\zeta_{2}\ell)\mathsf{E}\left[\max_{\exp(\ell)\leq k\leq\exp(\ell+1)}\left\|\sum_{t=1}^{k}\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)\left(\widehat{\mathbf{g}}_{t}-\mathbf{H}^{\top}\mathbf{g}_{t}\right)^{\top}\right\|\right]$
$\leq$ $\,\frac{1}{x}\sum_{\ell=0}^{\lfloor\log(T)\rfloor}\exp(-\zeta_{2}\ell)\frac{\exp(\ell+1)}{\min\{N,\exp(\ell+1)\}},$ (A.27)
where the last passage follows from Lemma A.7. If $N\geq T$, the conclusion follows trivially. On the other hand, if $\min\{N,\exp(\ell+1)\}=N$ for some $\ell$, we have $N=T^{\beta}$ with some $\beta\in(1/2+\epsilon_{\circ},1)$ under Assumption 6. Then, the RHS of (A.27) is bounded by
$\frac{1}{x}\sum_{\ell=0}^{\lfloor\log(T)\rfloor}\exp(-\zeta_{2}\ell)+\frac{\exp(-\beta\log(T))}{x}\sum_{\ell=0}^{\lfloor\log(T)\rfloor}\exp((1-\zeta_{2})\ell+1)\leq\frac{c_{0}}{x},$
provided that $1-\beta\leq\zeta_{2}$, which follows for $\zeta_{2}=1/2-\epsilon$ with some $\epsilon\in(0,\epsilon_{\circ})$. This proves the desired result. The proof of (A.26) takes analogous steps and we discuss it only briefly. Note that, setting $r=d=1$ and omitting $\mathbf{H}$ for simplicity,
$\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left|\sum_{t=1}^{k}g_{t}\left(\widehat{g}_{t}-g_{t}\right)\right|$
$\leq$ $\,\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left|\frac{1}{T}\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{g}_{s}g_{t}\gamma_{s,t}\right|+\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left|\frac{1}{T}\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{g}_{s}g_{t}\zeta_{s,t}\right|$
$\leq$ $\,+\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left|\frac{1}{T}\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{g}_{s}g_{t}\eta_{s,t}\right|+\max_{1\leq k\leq T}\frac{1}{k^{\zeta_{2}}}\left|\frac{1}{T}\sum_{t=1}^{k}\sum_{s=1}^{T}\widehat{g}_{s}g_{t}\xi_{s,t}\right|$ (A.28)
by applying (A.1) as in (A.4). Then, the proof proceeds as in the proof of (A.25) to each term in the RHS of (A.28) using Lemma A.8. ∎
Lemma A.11.
Suppose that Assumptions 16 hold with $\rho=1$ in Assumptions 1, 3 and 4. Then on a suitably enlarged probability space, there exists some constant $\zeta\in(0,1/2)$ such that
$\max_{1\leq k\leq T/2}\frac{1}{k^{\zeta}}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right.-$
$\qquad\qquad\qquad\left.\sum_{j=0}^{R}\mathbb{I}_{\{k_{j}\leq k\}}\cdot\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathbf{K}_{r}\mathbf{D}_{j}^{1/2}W_{1,dT}^{(j)}\left(\min(k,k_{j+1})-k_{j}\right)\right\|=O_{P}(1),$ (A.29)
$\max_{T/2<k<T}\frac{1}{(T-k)^{\zeta}}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right.-$
$\qquad\qquad\left.\sum_{j=0}^{R}\mathbb{I}_{\{k_{j+1}\geq k\}}\cdot\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathbf{K}_{r}\mathbf{D}_{j}^{1/2}W_{2,dT}^{(j)}\left(k_{j+1}-\max(k,k_{j})\right)\right\|=O_{P}(1),$ (A.30)
where $\mathbf{D}_{j}$ is defined in (A.24) and $W_{\ell,dT}^{(j)}(\cdot),\,\ell=1,2$, in Lemma A.9.
Proof.
The proof follows immediately from Lemmas A.9 and A.10. We prove (A.29) only since the arguments for (A.30) are analogous. Let $\zeta=\max(\zeta_{1},\zeta_{2})$ where $\zeta_{1}$ and $\zeta_{2}$ are defined in Lemmas A.9 and A.10, respectively. Then we have
$\max_{1\leq k\leq T/2}\frac{1}{k^{\zeta}}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\right.$
$\qquad\qquad\qquad\qquad\left.\sum_{j=0}^{R}\mathbb{I}_{\{k_{j}\leq k\}}\cdot\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathbf{K}_{r}\mathbf{D}_{j}^{1/2}W^{(j)}_{1,dT}\left(\min(k,k_{j+1})-k_{j}\right)\right\|$
$\leq$ $\,\max_{1\leq k\leq T/2}\frac{1}{k^{\zeta}}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}\right)\right\|$
$\,+\max_{1\leq k\leq T/2}\frac{1}{k^{\zeta}}\left\|\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathbf{K}_{r}\left(\sum_{t=1}^{k}\mathsf{Vech}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\right)-\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.\left.\sum_{j=0}^{R}\mathbb{I}_{\{k_{j}\leq k\}}\cdot\mathbf{D}_{j}^{1/2}W^{(j)}_{1,dT}\left(\min(k,k_{j+1})-k_{j}\right)\right)\right\|=:T_{1}+T_{2}.$
Lemma A.10 immediately yields that $T_{1}=O_{P}(1)$. We have $T_{2}$ bounded in light of Lemma A.9 since $\|\mathbf{H}\|=O_{P}(1)$ (see (A.11)) and $\|\mathbf{L}_{r}\|=O(1)$ and $\|\mathbf{K}_{r}\|=O(1)$ by their construction. ∎
The following two lemmas are useful in studying the behaviour of MOSUM statistics in (3) in the presence of multiple change points.
Lemma A.12.
Suppose that Assumptions 16 hold with $\rho=1$ in Assumptions 1, 3 and 4. Then it holds that
$\max_{0\leq k\leq T-\gamma}\frac{1}{\sqrt{\gamma}}\left\|\sum_{t=k+1}^{k+\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right\|=O_{P}(\sqrt{\log(T/\gamma)}).$
Proof.
By Lemma A.11, on a suitably enlarged probability space, it holds that
$\max_{0\leq k\leq T-\gamma}\frac{1}{\sqrt{\gamma}}\left\|\sum_{t=k+1}^{k+\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right\|$
$\leq$ $\frac{1}{\sqrt{\gamma}}\max_{0\leq k\leq T-\gamma}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\sum_{j=0}^{R}\mathbb{I}_{\{k_{j}\leq k\}}\cdot\mathbf{V}_{j}^{1/2}W^{(j)}_{1,dT}\left(\min(k,k_{j+1})-k_{j}\right)\right\|$
$+\frac{1}{\sqrt{\gamma}}\max_{0\leq k\leq T-\gamma}\left\|\sum_{t=1}^{k+\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\right.$
$\qquad\qquad\qquad\left.\sum_{j=0}^{R}\mathbb{I}_{\{k_{j}\leq k+\gamma\}}\cdot\mathbf{V}_{j}^{1/2}W^{(j)}_{1,dT}\left(\min(k+\gamma,k_{j+1})-k_{j}\right)\right\|$
$+\max_{0\leq k\leq T-\gamma}\frac{1}{\sqrt{\gamma}}\left\|\sum_{j=0}^{R}\mathbb{I}_{\{k<k_{j}\leq k+\gamma\}}\cdot\mathbf{V}_{j}^{1/2}W^{(j)}_{1,dT}\left(\min(k+\gamma,k_{j+1})-\max(k,k_{j})\right)\right\|$
=:absent:=:= : $\,T_{1}+T_{2}+T_{3}.$
Using Lemma A.11,
$T_{1}\leq$ $\,\frac{1}{\sqrt{\gamma}}\max_{0\leq k\leq T-\gamma}k^{\zeta}\cdot k^{-\zeta}\left\|\sum_{t=1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\right.$
$\qquad\qquad\qquad\qquad\qquad\left.\sum_{j=0}^{R}\mathbb{I}_{\{k_{j}\leq k\}}\cdot\mathbf{V}_{j}^{1/2}W^{(j)}_{1,dT}\left(\min(k,k_{j+1})-k_{j}\right)\right\|$
$=$ $\,\frac{1}{\sqrt{\gamma}}\cdot O_{P}(1)\max_{0\leq k\leq T-\gamma}k^{\zeta}=O_{P}\left(\frac{T^{\zeta}}{\sqrt{\gamma}}\right)=O_{P}\left(\frac{1}{\sqrt{\log(T/\gamma)}}\right),$
where the last equality follows from (8); the term $T_{2}$ is analogously bounded. From Theorem 1 in Shao, (1995) and the fact that there are finitely many change points, we have $T_{3}=O_{P}(\sqrt{\log(T)})$, which completes the proof. ∎
Lemma A.13.
Suppose that Assumptions 17 hold with $\rho=2$ in Assumptions 1, 3 and 4. Let us define $D_{T}=\min_{1\leq j\leq R}d_{j}\sqrt{\gamma}$. Then for any sequence $a_{T}$ satisfying $1\leq a_{T}\leq D_{T}$, and a (slowly varying) sequence $\omega_{T}\rightarrow\infty$, define
$\mathcal{M}_{T}^{(\ell)}=$$\,\left\{\max_{1\leq j\leq R}\max_{d_{j}^{-2}a_{T}\leq k\leq k_{j}-k_{j-1}}\frac{\sqrt{d_{j}^{-2}a_{T}}}{k}\left\|\sum_{t=k_{j}+\ell\gamma-k+1}^{k_{j}+\ell\gamma}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right\|_{F}\leq\omega_{T}\right\}$
$\bigcap\left\{\max_{1\leq j\leq R}\max_{d_{j}^{-2}a_{T}\leq k\leq k_{j}-k_{j-1}}\frac{\sqrt{d_{j}^{-2}a_{T}}}{k}\left\|\sum_{t=k_{j}+\ell\gamma+1}^{k_{j}+\ell\gamma+k}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right\|_{F}\leq\omega_{T}\right\}$
for some $\epsilon>0$ and $\ell\in\{0,\pm 1\}$. Then it holds that, as $\min(N,T)\rightarrow\infty$,
$\mathsf{P}\left(\cap_{\ell\in\{0,\pm 1\}}\mathcal{M}_{T}^{(\ell)}\right)\rightarrow 1.$ (A.31)
Proof.
We base the proof on Proposition 2.1 (c.ii) in Cho and Kirch, (2022), where a sufficient condition for (A.31) is that
$\mathsf{E}\left(\left\|\sum_{t=a+1}^{b}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\right\|^{2+\epsilon}\right)\leq c_{0}(b-a)^{1+\epsilon/2}$ (A.32)
for some $\epsilon>0$. This in turn follows if we show that
$\mathsf{E}\left(\left\|\sum_{t=a+1}^{b}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}\right)\right\|^{2+\epsilon}\right)$ $\leq c_{0}(b-a)^{1+\epsilon/2},$ (A.33)
$\mathsf{E}\left(\left\|\sum_{t=a+1}^{b}\mathsf{Vech}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\right)\right\|^{2+\epsilon}\right)$ $\leq c_{0}(b-a)^{1+\epsilon/2},$ (A.34)
together with (A.11). Equation (A.34) follows immediately from Proposition 4 of Berkes et al., (2011), which entails that
$\mathsf{E}\left(\left\|\sum_{t=a+1}^{b}\mathsf{Vech}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\right)\right\|^{2+\epsilon}\right)$
$=$ $\,\mathsf{E}\left(\left\|\sum_{j=0}^{R}\mathbb{I}_{\{a\leq k_{j}<b\}}\cdot\sum_{t=\max(a,k_{j})+1}^{\min(b,k_{j+1})}\mathsf{Vech}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}-\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\right)\right\|^{2+\epsilon}\right)$
$\leq$ $\,\sum_{j=0}^{R}\mathbb{I}_{\{a\leq k_{j}<b\}}\cdot\|\mathbf{A}_{j}\|^{4+2\epsilon}\mathsf{E}\left(\left\|\sum_{t=\max(a,k_{j})+1}^{\min(b,k_{j+1})}\mathsf{Vech}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\top}-\mathsf{E}\left(\mathbf{f}_{t}\mathbf{f}_{t}^{\top}\right)\right)\right\|^{2+\epsilon}\right)$
$\leq$ $\,c_{0}R(b-a)^{1+\epsilon/2},$
by Assumption 1, Assumption 5 (i) and (ii). As for (A.33), mechanically repeating the arguments in the proofs of Lemmas A.3 and A.4, we obtain that
$\mathsf{E}\left(\left\|\sum_{t=a+1}^{b}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}\right)\right\|^{2+\epsilon}\right)\leq c_{0}\left(\frac{b-a}{\min(N,b-a)}\right)^{2+\epsilon}.$
By elementary arguments
$\frac{b-a}{\min(N,b-a)}=\max\left(1,\frac{b-a}{N}\right)\leq\max\left\{1,\frac{T^{1/2+\epsilon^{\prime}}}{N}(b-a)^{1/2-\epsilon^{\prime}}\right\}=o\left((b-a)^{1/2-\epsilon^{\prime}}\right)$
for some $\epsilon^{\prime}\in(0,\epsilon_{\circ})$ under Assumption 6. Therefore,
$\mathsf{E}\left(\left\|\sum_{t=a+1}^{b}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}\right)\right\|^{2+\epsilon}\right)\leq c_{0}(b-a)^{1/2+\epsilon/2}.$
Putting all together, the condition in (A.32), which completes the proof. ∎
A.2 Proof of Theorem 1
Proof of Theorem 1 (a).
Let us define a symmetric, $d\times d$-matrix
$\widetilde{\mathbf{V}}=\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathbf{K}_{r}\mathbf{D}_{j}\mathbf{K}_{r}^{\top}(\mathbf{H}\otimes\mathbf{H})\mathbf{L}_{r}^{\top}.$
From Lemma A.6 and its proof, we have $\mathbf{H}$ asymptotically invertible and $\|\mathbf{H}\|=O_{P}(1)$. Also, from that $\|\mathbf{D}\|=O(1)$ (due to Assumption 1 (i)), $\|\mathbf{D}^{-1}\|=O(1)$ (Assumption 1 (iii)) and
$\Lambda_{\min}(\widetilde{\mathbf{V}})\geq\Lambda_{\min}(\mathbf{D})\left\|\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathbf{K}_{r}\right\|_{F}^{2},$
we have $\widetilde{\mathbf{V}}$ asymptotically invertible with
$\|\widetilde{\mathbf{V}}\|=O_{P}(1)\text{ \ and \ }\|\widetilde{\mathbf{V}}^{-1}\|=O_{P}(1).$ (A.35)
Then by Lemma A.11 (with $R=0$ under $\mathcal{H}_{0}$), there exist two independent $d$-dimensional Wiener processes $W_{\ell,dT}(\cdot),\,\ell=1,2$, such that
$\max_{1\leq k\leq T}\frac{1}{k^{\zeta}}\left\|\sum_{t=1}^{k}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-W_{dT}(k)\right\|$ $=O_{P}(1),$ (A.36)
where $W_{dT}(k)=W_{1,dT}(\min(k,T/2))+(W_{2,dT}(T/2)-W_{2,dT}(T-k))\cdot\mathbb{I}_{\{k>T/2\}}$, and $\zeta=\max(\zeta_{1},\zeta_{2})$ with $\zeta_{1}$ and $\zeta_{2}$ defined in Lemmas A.9 and A.10, respectively. Following Theorem S2.1 in Berkes et al., (2014), which is referred to in the proof of Lemma A.9, we have $\zeta_{1}=2/\nu$ with $\nu$ denote the largest number such that $\mathsf{E}(|g_{t}|^{\nu})<\infty$; under Assumption 1 (i), we can set e.g. $\nu=8$. Further, inspecting the proof of Lemma A.10, it emerges that whenever
$\frac{1}{2}+\epsilon_{\circ}<\beta=\frac{\log(N)}{\log(T)}\leq 1,$
it must hold that $1-\zeta_{2}\leq\beta$, whereas $\zeta_{2}>0$ can be arbitrarily small when $\beta>1$. Hence we set $\zeta_{2}=1-\min(1,\beta)$. Thus, the statement in (A.36) holds with $\zeta$ chosen as in (7).
The rest of the proof now is similar to that of Theorem 2.1 in Hušková and Slabỳ, (2001). Note that
$\sum_{t=k+1}^{k+\gamma}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)\\ =\sum_{t=1}^{k+\gamma}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\sum_{t=1}^{k}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right).$
It holds that
$\max_{1\leq k\leq T-\gamma}\frac{1}{\sqrt{2\gamma}}\left\|\sum_{t=1}^{k+\gamma}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-W_{dT}(k+\gamma)\right\|$
$=\frac{1}{\sqrt{2\gamma}}\max_{1\leq k\leq T-\gamma}\frac{1}{(k+\gamma)^{\zeta}}\left\|\sum_{t=1}^{k+\gamma}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-W_{dT}(k+\gamma)\right\|$
$\qquad\qquad\qquad\cdot\max_{1\leq k\leq T-\gamma}(k+\gamma)^{\zeta}=O_{P}\left(T^{\zeta}\gamma^{-1/2}\right)=o_{P}\left(\log^{-1/2}(T/\gamma)\right)$
under (A.36), (7) and (8). Similarly,
$\max_{\gamma<k\leq T}\frac{1}{\sqrt{2\gamma}}\left\|\sum_{t=1}^{k}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-W_{dT}(k)\right\|$
$=$ $\,O_{P}\left(T^{\zeta}\gamma^{-1/2}\right)=o_{P}\left(\log^{-1/2}(T/\gamma)\right).$
From the above, we obtain
$\frac{1}{\sqrt{2\gamma}}\max_{\gamma\leq k\leq T-\gamma}\left\|\sum_{t=k+1}^{k+\gamma}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\left(W_{dT}(k+\gamma)-W_{dT}(k)\right)\right\|\\ =o_{P}\left(\log^{-1/2}(T/\gamma)\right).$ (A.37)
Analogously, we can show that
$\frac{1}{\sqrt{2\gamma}}\max_{\gamma\leq k\leq T-\gamma}\left\|\sum_{t=k-\gamma+1}^{k}\widetilde{\mathbf{V}}^{-1/2}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathbf{H}\right)-\left(W_{dT}(k)-W_{dT}(k-\gamma)\right)\right\|\\ =o_{P}\left(\log^{-1/2}(T/\gamma)\right).$ (A.38)
Combining (A.37) and (A.38), and from the fact that $\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})$ is time-invariant, we obtain
$\max_{\gamma\leq k\leq T-\gamma}\left\|\widetilde{\mathbf{V}}^{-1/2}\mathbf{M}_{N,T,\gamma}(k)\right\|\\ =\frac{1}{\sqrt{2\gamma}}\max_{\gamma\leq k\leq T-\gamma}\left\|W_{dT}(k+\gamma)-W_{dT}(k-\gamma)\right\|+o_{P}\left(\log^{-1/2}(T/\gamma)\right).$ (A.39)
Let $k=\lfloor\gamma t\rfloor$ with $11\leq t\leq T/\gamma-$ 1. On account of (A.39), we will study
$\frac{1}{\sqrt{2\gamma}}\max_{\gamma\leq k\leq T-\gamma}\left\|W_{dT}(k+\gamma)-W_{dT}(k-\gamma)\right\|$
$=$ $\,\frac{1}{\sqrt{2\gamma}}\max_{1\leq t\leq T/\gamma}\left\|W_{dT}\left(\lfloor\gamma t\rfloor+\gamma\right)-W_{dT}\left(\lfloor\gamma t\rfloor-\gamma\right)\right\|$
=$\overset{\mathcal{D}}{=}$ $\,\frac{1}{\sqrt{2}}\max_{1\leq t\leq T/\gamma-1}\left\|W_{dT}(t+1)-W_{dT}(t-1)\right\|,$
having used the scale transformation of the Wiener process. Since the distribution of $W_{dT}(t)$ does not depend on $T$, we also have that, as $T\rightarrow\infty$ with (8),
$\frac{1}{\sqrt{2}}\max_{1\leq t\leq T/\gamma-1}\left\|W_{dT}(t+1)-W_{dT}(t-1)\right\|\rightarrow\frac{1}{\sqrt{2}}\max_{1\leq t<\infty}\left\|W_{dT}(t+1)-W_{dT}(t-1)\right\|$
almost surely. The $d$-dimensional process
$\omega(t)=\frac{1}{\sqrt{2}}\left(W_{dT}(t+1)-W_{dT}(t-1)\right),$ (A.40)
[…]
$\mathsf{E}\left(\omega_{i}(t)\omega_{i}(t+h)\right)=\left\{\begin{array}[]{ll}1-\frac{1}{2}|h|&\text{for \ }0\leq|h|\leq 1,\\ 0&\text{for \ }|h|>2.\end{array}\right.$ (A.43)
[…]
$\lim_{\min(N,T)\rightarrow\infty}\mathsf{P}\left(a\left(\frac{T}{\gamma}\right)\max_{\gamma\leq k\leq T-\gamma}\left|\mathbf{M}_{N,T,\gamma}^{\top}(k)\widetilde{\mathbf{V}}^{-1}\mathbf{M}_{N,T,\gamma}(k)\right|^{1/2}-b_{d}\left(\frac{T}{\gamma}\right)\leq x\right)$
$=\exp\left(-2\exp(-x)\right).$ (A.44)
This, together with (A.35), implies that
$\max_{\gamma\leq k\leq T-\gamma}\left\|\mathbf{M}_{N,T,\gamma}(k)\right\|=O_{P}\left(\sqrt{\log(T/\gamma)}\right).$ (A.45)
Also, we have
$\left\|\widetilde{\mathbf{V}}-\mathbf{V}\right\|$ $=O_{P}\left(\left\|(\mathbf{H}\otimes\mathbf{H})-(\mathbf{H}_{0}\otimes\mathbf{H}_{0})\right\|\right)=O_{P}\left(\left\|\mathbf{H}-\mathbf{H}_{0}\right\|\right)$
$=O_{P}\left(C_{NT}^{-1}\right)=o_{P}\left(\log^{-1}(T/\gamma)\right)$ (A.46)
by Assumption 6 and Lemma A.6. Then, e.g. by Lemma 4.1 of Powers and Størmer, (1970), we have
$\left\|\widetilde{\mathbf{V}}^{-1/2}-\mathbf{V}^{-1/2}\right\|$ $=o_{P}\left(\log^{-1}(T/\gamma)\right).$
This, together with (A.45), establishes that we can replace $\widetilde{\mathbf{V}}$ with $\mathbf{V}$ and continue to have the asymptotic distribution in (A.44) hold, since
$\left|\max_{\gamma\leq k\leq T-\gamma}\left\|\widetilde{\mathbf{V}}^{-1/2}\mathbf{M}_{N,T,\gamma}(k)\right\|-\max_{\gamma\leq k\leq T-\gamma}\left\|\mathbf{V}^{-1/2}\mathbf{M}_{N,T,\gamma}(k)\right\|\right|$
$\leq$ $\,\max_{\gamma\leq k\leq T-\gamma}\left\|\left(\widetilde{\mathbf{V}}^{-1/2}-\mathbf{V}^{-1/2}\right)\mathbf{M}_{N,T,\gamma}(k)\right\|$
$\leq$ $\,\left\|\widetilde{\mathbf{V}}^{-1/2}-\mathbf{V}^{-1/2}\right\|\cdot\max_{\gamma\leq k\leq T-\gamma}\left\|\mathbf{M}_{N,T,\gamma}(k)\right\|$
$=$ $\,o_{P}\left(\log^{-1}(T/\gamma)\right)\cdot O_{P}\left(\sqrt{\log(T/\gamma)}\right)=o_{P}\left(a^{-1}(T/\gamma)\right).$ (A.47)
Proof of Theorem 1 (b).
Arguments analogous to those leading to (A.47) can be adopted under (11), and thus we omit the proof. ∎
A.3 Proof of Proposition 2
The proof follows similar passages to Han and Inoue, (2015), and therefore we focus only on some aspects of it. We will use the following notations
$\widehat{\bm{\Gamma}}(\ell)$ $=\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t-\ell}^{\top}\text{ \ with \ }\mathbf{Z}_{t}=\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{I}_{r}\right),$
$\bm{\Gamma}(\ell)$ $=\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\text{ \ with \ }\widetilde{\mathbf{U}}_{t}=\mathsf{Vech}\left(\mathbf{H}_{0}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}_{0}-\mathbf{I}_{r}\right),\text{ \ and}$
$\mathbf{U}_{t}$ $=\mathsf{Vech}\left(\mathbf{H}^{\top}\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\mathbf{H}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right).$
Noting that $\mathbf{H}_{0}^{\top}\bm{\Sigma}_{G}\mathbf{H}_{0}=\mathbf{I}_{r}$ (see the proof of Lemma A.6), it follows that
$\mathbf{V}=\bm{\Gamma}(0)+\sum_{\ell=1}^{\infty}\left(\bm{\Gamma}(\ell)+\bm{\Gamma}(\ell)^{\top}\right),$
so that
$\widehat{\mathbf{V}}-\mathbf{V}=$ $\,\left(\widehat{\bm{\Gamma}}(0)-\bm{\Gamma}(0)\right)+\sum_{\ell=1}^{m}\left(1-\frac{\ell}{m+1}\right)\left[\left(\widehat{\bm{\Gamma}}(\ell)-\bm{\Gamma}(\ell)\right)+\left(\widehat{\bm{\Gamma}}(\ell)-\bm{\Gamma}(\ell)\right)^{\top}\right]$
$+\sum_{\ell=1}^{m}\frac{\ell}{m+1}\left(\bm{\Gamma}(\ell)+\bm{\Gamma}(\ell)^{\top}\right)+\sum_{\ell=m+1}^{\infty}\left(\bm{\Gamma}(\ell)+\bm{\Gamma}(\ell)^{\top}\right).$ (A.48)
Since $\mathbf{g}_{t}$ is an $\mathcal{L}_{8+\epsilon}$-decomposable Bernoulli shift with $a>2$, it is easy to see (cfr. the proof of Lemma A.9) that $\mathsf{Vech}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})$ is an $\mathcal{L}_{4+\epsilon/2}$-decomposable Bernoulli shift, also with $a>2$. The covariance summability of Bernoulli shifts (see e.g. Lemma D.4 in Horváth and Trapani,, 2023) entails that
$\left\|\sum_{\ell=1}^{m}\frac{\ell}{m+1}\left(\bm{\Gamma}(\ell)+\bm{\Gamma}(\ell)^{\top}\right)\right\|$ $=O\left(\frac{1}{m}\right),$ (A.49)
$\left\|\sum_{\ell=m+1}^{\infty}\left(\bm{\Gamma}(\ell)+\bm{\Gamma}(\ell)^{\top}\right)\right\|$ $=O\left(\frac{1}{m}\right).$ (A.50)
We now bound the rest of the terms in (A.48). First, note that
$\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t-\ell}^{\top}=$ $\,\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{U}_{t}\mathbf{U}_{t-\ell}^{\top}+\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{U}_{t}(\mathbf{Z}_{t-\ell}-\mathbf{U}_{t-\ell})^{\top}$
$+\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}(\mathbf{Z}_{t}-\mathbf{U}_{t})\mathbf{U}_{t-\ell}^{\top}+\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}(\mathbf{Z}_{t}-\mathbf{U}_{t})(\mathbf{Z}_{t-\ell}-\mathbf{U}_{t-\ell})^{\top}$
=:absent:=:= : $\,T_{1}+T_{2}+T_{3}+T_{4}.$ (A.51)
First we study $T_{2}$ in (A.51). For simplicity, let $r=d=1$ and omit $\mathbf{H}$ noting that $\|\mathbf{H}\|=O_{P}(1)$ due to (A.11). Further, we may treat $\mathsf{E}(g_{t}^{2})=1$. Then, we can write
$T_{2}$ $=\frac{1}{T}\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)\left(\sum_{\ell=1}^{m}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)^{2}+2\sum_{\ell=1}^{m}g_{t-\ell}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)\right)\mathbb{I}_{\{\ell<t\}}$
$\leq\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)^{2}\right)^{2}+4\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}g_{t-\ell}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)\right)^{2}\right)^{1/2}$
$\leq\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)^{2}\right)^{2}\right)^{1/2}$
$\quad+\frac{2}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}g_{t-\ell}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)\right)^{2}\right)^{1/2}=:T_{2,1}+T_{2,2}.$ (A.52)
Assumption 1 (i) entails $\sum_{t=1}^{T}(g_{t}^{2}-\mathsf{E}(g_{t}^{2}))^{2}=O_{P}(T)$. Also, by (A.1),
$\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)^{2}\right)^{2}$
$\leq$ $\,\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|^{2}\right)^{2}+\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|^{2}\right)^{2}$
$\,+\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t-\ell}\right|^{2}\right)^{2}+\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\xi_{s,t-\ell}\right|^{2}\right)^{2}.$
First, we have
$\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|^{2}\right)^{2}$
$\leq$ $\,T^{-4}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left|\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|^{4}\leq T^{-4}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left(\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right)^{2}\left(\sum_{s=1}^{T}\gamma_{s,t-\ell}^{2}\right)^{2}$
$\leq$ $\,T^{-2}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left(\sum_{s=1}^{T}\gamma_{s,t-\ell}^{2}\right)^{2}\leq T^{-2}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left(\sum_{s=1}^{T}|\gamma_{s,t-\ell}|\right)^{4}\leq c_{0}T^{-1}m^{2}$
by Assumption 3 (ii). Thus, it holds that
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|^{2}\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{T}\right).$
Using similar arguments,
$\mathsf{E}\left[\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|^{2}\right)^{2}\right]$
$\leq$ $\,\mathsf{E}\left[T^{-2}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left(\sum_{s=1}^{T}\zeta_{s,t-\ell}^{2}\right)^{2}\right]\leq T^{-1}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathsf{E}\left(\zeta_{s,t-\ell}^{4}\right)\leq c_{0}\frac{m^{2}T}{N^{2}}$
by Assumption 3 (iii), whereby
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|^{2}\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{N}\right).$
Similarly we get
$\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t-\ell}\right|^{2}\right)^{2}$
$\leq$ $\,T^{-1}m\sum_{\ell=1}^{m}\sum_{t=1}^{T}\sum_{s=1}^{T}\left|\eta_{s,t-\ell}\right|^{4}=T^{-1}m\left(\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t-\ell}\right|^{4}\right)\left(\sum_{s=1}^{T}\left|g_{s}\right|^{4}\right)$
$=$ $\,O_{P}\left(T\right)T^{-1}m\left(\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t-\ell}\right|^{4}\right)=O_{P}\left(\frac{m^{2}T}{N^{2}}\right),$
having used Assumption 3 (iv) in the final passage. Thus
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t-\ell}\right|^{2}\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{N}\right).$
By the same arguments, we also obtain
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}\sum_{s=1}^{T}\widehat{g}_{s}\xi_{s,t-\ell}\right|^{2}\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{N}\right),$
and therefore in (A.52), $T_{2,1}=O_{P}(mC_{NT}^{-2})$. Similarly, we note that
$\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}g_{t-\ell}\left(\widehat{g}_{t-\ell}-g_{t-\ell}\right)\right)^{2}$
$\leq$ $\,\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|\right)^{2}+\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|\right)^{2}$
$\,+\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t-\ell}\right|\right)^{2}+\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\xi_{s,t-\ell}\right|\right)^{2}.$
From Assumption 3 (ii), it holds that
$\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|\right)^{2}$
$\leq$ $\,mT^{-2}\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\left|\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|^{2}\leq mT^{-2}\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\left(\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right)\left(\sum_{s=1}^{T}\gamma_{s,t-\ell}^{2}\right)$
$\leq$ $\,mT^{-1}\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\left(\sum_{s=1}^{T}\left|\gamma_{s,t-\ell}\right|\right)^{2}=O_{P}(m^{2}),$
with which we obtain
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\gamma_{s,t-\ell}\right|\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{T^{1/2}}\right).$
Similarly,
$\mathsf{E}\left[\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|\right)^{2}\right]$
$\leq$ $\,mT^{-2}\mathsf{E}\left[\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\left|\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|^{2}\right]\leq mT^{-2}\mathsf{E}\left[\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\left(\sum_{s=1}^{T}\widehat{g}_{s}^{2}\right)\left(\sum_{s=1}^{T}\zeta_{s,t-\ell}^{2}\right)\right]$
$\leq$ $\,mT^{-1}\left(\sum_{\ell=1}^{m}\sum_{t=1}^{T}\mathsf{E}\left(|g_{t-\ell}|^{4}\right)\right)^{1/2}\left(\mathsf{E}\left[\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left(\sum_{s=1}^{T}\zeta_{s,t-\ell}^{2}\right)^{2}\right]\right)^{1/2}$
$\leq$ $\,mT^{-1}O\left(\left(mT\right)^{1/2}\right)\left(T\sum_{\ell=1}^{m}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathsf{E}\left(\left|\zeta_{s,t-\ell}\right|^{4}\right)\right)^{1/2}=O\left(\frac{m^{2}T}{N}\right),$
by Assumption 3 (ii), so that
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\zeta_{s,t-\ell}\right|\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{N^{1/2}}\right).$
Analogously,
$\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t-\ell}\right|\right)^{2}$
$\leq$ $\,mT^{-1}\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\sum_{s=1}^{T}\left|g_{s}\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t-\ell}\right|^{2}=mT^{-1}\left(\sum_{s=1}^{T}g_{s}^{2}\right)\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{2}\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t-\ell}\right|^{2}$
$\leq$ $\,mT^{-1}O_{P}(T)\left(\sum_{\ell=1}^{m}\sum_{t=1}^{T}g_{t-\ell}^{4}\right)^{1/2}\left(\sum_{\ell=1}^{m}\sum_{t=1}^{T}\left|\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}e_{i,t-\ell}\right|^{4}\right)^{1/2}$
$=$ $\,mT^{-1}O_{P}\left(T\cdot\left(mT\right)^{1/2}\cdot\frac{\left(mT\right)^{1/2}}{N}\right)=O_{P}\left(\frac{m^{2}T}{N}\right).$
Hence,
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\eta_{s,t-\ell}\right|\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{N^{1/2}}\right).$
The same passages, in essence, yield
$\frac{1}{T}\left(\sum_{t=1}^{T}\left(g_{t}^{2}-\mathsf{E}(g_{t}^{2})\right)^{2}\right)^{1/2}\left(\sum_{t=1}^{T}\left(\sum_{\ell=1}^{m}\left|\frac{1}{T}g_{t-\ell}\sum_{s=1}^{T}\widehat{g}_{s}\xi_{s,t-\ell}\right|\right)^{2}\right)^{1/2}=O_{P}\left(\frac{m}{N^{1/2}}\right),$
so that $T_{2,2}=O_{P}(mC_{NT}^{-1})$ in (A.52). Therefore, we finally have
$T_{2}=O_{P}\left(\frac{m}{C_{NT}}\right).$
Following the analogous arguments, $T_{3}$ and $T_{4}$ are similarly bounded, from which we conclude that
$\left\|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{U}_{t}\mathbf{U}_{t-\ell}^{\top}\right\|=O_{P}\left(\frac{m}{C_{NT}}\right),$ (A.53)
and repeating essentially the same passages, it can be shown that
$\left\|\frac{1}{T}\sum_{t=1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}-\frac{1}{T}\sum_{t=1}^{T}\mathbf{U}_{t}\mathbf{U}_{t}^{\top}\right\|=O_{P}\left(\frac{1}{C_{NT}}\right).$
Next, note that by definition, we can write
$\mathbf{U}_{t}\mathbf{U}_{t-\ell}^{\top}=\mathbf{L}_{r}(\mathbf{H}^{\top}\otimes\mathbf{H}^{\top})\mathsf{Vec}\left(\mathbf{g}_{t}\mathbf{g}_{t}^{\top}\right)\mathsf{Vec}\left(\mathbf{g}_{t-\ell}\mathbf{g}_{t-\ell}^{\top}\right)^{\top}(\mathbf{H}\otimes\mathbf{H})\mathbf{L}_{r}^{\top},$
and a similar representation holds for $\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}$ with $\mathbf{H}_{0}$ replacing $\mathbf{H}$. Then, by Lemma A.6, it can be verified that
$\left\|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{U}_{t}\mathbf{U}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right\|=O_{P}\left(\frac{m}{C_{NT}}\right).$ (A.54)
We now consider bounding
$\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)$
and again, we set $r=d=1$ for simplicity. Under Assumption 1 (i), it is easy to see that for all $j\geq 0$, the sequence $S_{t,\ell}=\widetilde{U}_{t}\widetilde{U}_{t-\ell}^{\top}-\mathsf{E}(\widetilde{U}_{t}\widetilde{U}_{t-\ell}^{\top})$ is an $\mathcal{L}_{2}$-decomposable Bernoulli shift with $a>2$. Hence
$\mathsf{E}\left(\left|\frac{1}{T}\sum_{\ell=1}^{m}\sum_{t=\ell+1}^{T}S_{t,\ell}\right|^{2}\right)\leq T^{-2}m\sum_{\ell=1}^{m}\mathsf{E}\left(\left|\sum_{t=\ell+1}^{T}S_{t,\ell}\right|^{2}\right)\leq\frac{c_{0}m^{2}}{T},$
where the last passage follows from Proposition 4 in Berkes et al., (2011). Hence we have
$\left|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\right|=O_{P}\left(\frac{m}{T^{1/2}}\right).$ (A.55)
Thus, putting together (A.53), (A.54) and (A.55), it follows that
$\left|\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t-\ell}^{\top}-\sum_{\ell=1}^{m}\frac{1}{T}\sum_{t=\ell+1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t-\ell}^{\top}\right)\right|=O_{P}\left(\frac{m}{C_{NT}}\right)$ (A.56)
and similarly,
$\left|\frac{1}{T}\sum_{t=1}^{T}\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}-\frac{1}{T}\sum_{t=1}^{T}\mathsf{E}\left(\widetilde{\mathbf{U}}_{t}\widetilde{\mathbf{U}}_{t}^{\top}\right)\right|=O_{P}\left(\frac{1}{C_{NT}}\right).$ (A.57)
Finally, using (A.49), (A.50), (A.56) and (A.57) in (A.48), it yields
$\left|\widehat{\mathbf{V}}-\mathbf{V}\right|=O_{P}\left(\frac{m}{C_{NT}}\right)+O\left(\frac{1}{m}\right)=O_{P}\left(\frac{1}{\log(T/\gamma)}\right),$
from the conditions made in (13) on $m$.
A.4 Proof of Theorem 3
WLOG, we may regard $\mathbf{V}=\mathbf{I}_{r}$, which does not alter the arguments as $\mathbf{V}$ is a positive definite matrix under Assumption 1 (iii) with bounded eigenvalues. Also for simplicity, we write $\mathbf{M}(k)=\mathbf{M}_{N,T,\gamma}(k)$ and $\mathcal{T}(k)=\mathcal{T}_{N,T,\gamma}(k)$. Decompose $\mathbf{M}(k)$ as
$\mathbf{M}(k)$ $=\frac{1}{\sqrt{2\gamma}}\left[\sum_{t=k+1}^{k+\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)-\sum_{t=k-\gamma+1}^{k}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)\right]$
$\qquad+\frac{1}{\sqrt{2\gamma}}\left[\sum_{t=k+1}^{k+\gamma}\mathsf{Vech}\left(\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)-\sum_{t=k-\gamma+1}^{k}\mathsf{Vech}\left(\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)\right]$
$=:\mathbf{N}(k)+\mathbf{S}(k).$
Then, we can write
$\frac{1}{2}\left(\left\|\mathbf{S}(k)\right\|^{2}+\left\|\mathbf{N}(k)\right\|^{2}\right)\leq(\mathcal{T}(k))^{2}=\left\|\mathbf{S}(k)+\mathbf{N}(k)\right\|^{2}\leq 2\left(\left\|\mathbf{S}(k)\right\|^{2}+\left\|\mathbf{N}(k)\right\|^{2}\right).$ (A.58)
From Lemma A.12,
$\max_{\gamma\leq k\leq T-\gamma}\|\mathbf{N}(k)\|=O_{P}\left(\sqrt{\log(T/\gamma)}\right).$ (A.59)
By definition of $\bm{\delta}_{j}$, we have
$\mathbf{S}(k)=\left\{\begin{array}[]{ll}\frac{\gamma-|k-k_{j}|}{\sqrt{2\gamma}}\;\mathsf{Vech}\left(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H}\right)&\text{if \ }k_{j}-\gamma+1\leq k\leq k_{j}+\gamma-1,\\ \mathbf{0}&\text{if \ }\min_{1\leq j\leq R}|k-k_{j}|\geq\gamma.\end{array}\right.$ (A.60)
Further, thanks to Lemma A.6, there exists some event $\mathcal{H}_{N,T}$ satisfying $\mathsf{P}(\mathcal{H}_{N,T})\rightarrow 1$ as $\min(N,T)\rightarrow\infty$ such that on $\mathcal{H}_{N,T}$,
$\left\|\mathsf{Vech}\left(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H}\right)\right\|$ $\geq\frac{1}{2}\left\|\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H}\right\|\geq\frac{1}{2}\Lambda_{\min}(\mathbf{H}^{\top}\mathbf{H})\|\bm{\delta}_{j}\|$
$\geq\frac{1}{2}\left(1-\left\|\mathbf{H}-\mathbf{H}_{0}\right\|\right)\|\bm{\delta}_{j}\|\geq\frac{1}{4}d_{j},$ (A.61)
and similarly
$\left\|\mathsf{Vech}\left(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H}\right)\right\|\leq\Lambda_{\max}(\mathbf{H}^{\top}\mathbf{H})\|\bm{\delta}_{j}\|\leq\frac{3}{2}d_{j}.$ (A.62)
Proof of Theorem 3 (a).
Consider for $j=1,\ldots,R$,
$\mathcal{S}_{T,j}=\left\{\mathcal{T}(k_{j})\geq\max\left(\max_{k:\,|k-k_{j}|>(1-\eta)\gamma}\mathcal{T}(k),\,\widetilde{D}_{T,\gamma}(\alpha)\cdot\omega_{T}^{(1)}\right)\right\}\cap\mathcal{H}_{N,T}$
and $\mathcal{S}_{T}=\bigcap_{1\leq j\leq R}\mathcal{S}_{T,j}$. Then for any $\alpha,\eta\in(0,1)$, we have
$(\mathcal{T}(k_{j}))^{2}$ $\geq\frac{1}{16}d_{j}^{2}\gamma+O_{P}(\log(T/\gamma))=\frac{1}{16}d_{j}^{2}\gamma(1+o_{P}(1)),$
$\max_{\begin{subarray}{c}k:\,|k-k_{j}|>(1-\eta)\gamma\\ 1\leq j\leq R\end{subarray}}(\mathcal{T}(k))^{2}$ $\leq\eta^{2}(\mathcal{T}(k_{j}))^{2}+O_{P}(\log(T/\gamma))=\eta^{2}(\mathcal{T}(k_{j}))^{2}(1+o_{P}(1))$
under Assumption 7, by (A.58), (A.59), (A.60) and (A.61), where the $O_{P}$-bounds hold uniformly over $j$ and $k$. Combined with that $\widetilde{D}_{T,\gamma}(\alpha)\asymp\sqrt{\log(T/\gamma)}$ for any $\alpha\in(0,1)$, we have $\mathsf{P}(\mathcal{S}_{T})\rightarrow 1$ as $\min(N,T)\rightarrow\infty$. Also defining
$\widetilde{\mathcal{S}}_{T,j}=\bigcap_{0\leq q\leq\lfloor 2/\eta\rfloor-2}\left[\left\{\mathcal{T}(k_{j}+q\eta\gamma/2)\geq\max_{k_{j}+(q+1)\eta\gamma/2\leq k\leq k_{j}+\gamma}\mathcal{T}(k)\right\}\right.\\ \left.\bigcap\left\{\mathcal{T}(k_{j}-q\eta\gamma/2)\geq\max_{k_{j}-\gamma\leq k\leq k_{j}-(q+1)\eta\gamma/2}\mathcal{T}(k)\right\}\right]\cap\mathcal{H}_{N,T},$
by the analogous arguments, we have $\mathsf{P}(\widetilde{\mathcal{S}}_{T,j})\rightarrow 1$ and hence $\mathsf{P}(\widetilde{\mathcal{S}}_{T})\rightarrow 1$ where $\widetilde{\mathcal{S}}_{T}=\cap_{j=1}^{R}\widetilde{\mathcal{S}}_{T,j}$. On $\mathcal{S}_{T}\cap\widetilde{\mathcal{S}}_{T}$, we detect exactly one change point estimator within the radius of $\eta\gamma/2$ for each change point according to the rule (4). Further, due to (A.59) and (A.60),
$\mathsf{P}\left\{\max_{\begin{subarray}{c}k:\,|k-k_{j}|\geq\gamma\\ 1\leq j\leq R\end{subarray}}\mathcal{T}(k)>\widetilde{D}_{T,\gamma}(\alpha)\cdot\omega_{T}^{(1)}\right\}\rightarrow 0\text{ \ as \ }\min(N,T)\rightarrow\infty,$
which guarantees that no estimator is detected outside the radius of $\gamma$ from each change point. Altogether, the above arguments show that
$\mathsf{P}\left(\widehat{R}=R;\,\max_{1\leq j\leq R}|\widehat{k}_{j}-k_{j}|\leq\eta\gamma/2\right)\rightarrow 1\text{ \ as \ }\min(N,T)\rightarrow\infty.$
Proof of Theorem 3 (b).
For each $j$, recall that $|\widehat{k}_{j}-k_{j}|\leq\gamma$, on $\mathcal{S}_{T}\cap\widetilde{\mathcal{S}}_{T}$. WLOG, suppose that $\widehat{k}_{j}\leq k_{j}$ and define $\widetilde{\mathcal{T}}_{j}(k)=(\mathcal{T}(k))^{2}-(\mathcal{T}(k_{j}))^{2}$. Then, recalling $\omega_{T}$ defined in Lemma A.13, let us consider
$\left\{Cd_{j}^{2}(\widehat{k}_{j}-k_{j})\leq-\omega_{T}^{2}\right\}\subset\left\{\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\widetilde{\mathcal{T}}_{j}(k)\geq\max_{k_{j}-Cd_{j}^{2}\omega_{T}^{2}+1\leq k\leq k_{j}+\gamma}\widetilde{\mathcal{T}}_{j}(k)\right\}\\ \subset\left\{\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\widetilde{\mathcal{T}}_{j}(k)\geq 0\right\}$
for some fixed $C\in(0,\infty)$. We can decompose $\widetilde{\mathcal{T}}_{j}(k)$ as
$\widetilde{\mathcal{T}}_{j}(k)$ $=\left(\mathbf{N}(k)-\mathbf{N}(k_{j})+\mathbf{S}(k)-\mathbf{S}(k_{j})\right)^{\top}\left(\mathbf{N}(k)+\mathbf{N}(k_{j})+\mathbf{S}(k)+\mathbf{S}(k_{j})\right)$
$=-\left(\mathbf{S}(k_{j})-\mathbf{S}(k)\right)^{\top}\left(\mathbf{S}(k_{j})+\mathbf{S}(k)\right)+\left(\mathbf{N}(k)-\mathbf{N}(k_{j})\right)\left(\mathbf{S}(k)+\mathbf{S}(k_{j})\right)$
$+\left(\mathbf{N}(k)+\mathbf{N}(k_{j})\right)\left(\mathbf{S}(k)-\mathbf{S}(k_{j})\right)+\left(\mathbf{N}(k)-\mathbf{N}(k_{j})\right)\left(\mathbf{N}(k_{j})+\mathbf{N}(k)\right)$
$=:\widetilde{\mathcal{T}}_{j,1}(k)+\widetilde{\mathcal{T}}_{j,2}(k)+\widetilde{\mathcal{T}}_{j,3}(k)+\widetilde{\mathcal{T}}_{j,4}(k).$
From (A.60), (A.61) and (A.62), we have $\widetilde{\mathcal{T}}_{j,1}(k)<0$ and
$\left|\widetilde{\mathcal{T}}_{j,1}(k)\right|=\frac{(2\gamma-|k-k_{j}|)|k-k_{j}|}{2\gamma}\left\|\mathsf{Vech}(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H})\right\|^{2}\geq\frac{1}{2}\left\|\mathsf{Vech}(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H})\right\|^{2}|k-k_{j}|,$
$\left\|\mathbf{S}(k_{j})-\mathbf{S}(k)\right\|=\frac{|k-k_{j}|}{\sqrt{2\gamma}}\left\|\mathsf{Vech}(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H})\right\|,$
$\left\|\mathbf{S}(k_{j})+\mathbf{S}(k)\right\|\leq\sqrt{2\gamma}\left\|\mathsf{Vech}(\mathbf{H}^{\top}\bm{\delta}_{j}\mathbf{H})\right\|.$ (A.63)
Then, using the arguments from the proof of Theorem 3.2 in Eichinger and Kirch, (2018),
$\mathsf{P}\left(\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\widetilde{T}_{j}(k)\geq 0,\,\mathcal{S}_{T}\cap\widetilde{\mathcal{S}}_{T}\right)$
$\leq$ $\,\mathsf{P}\left(\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\left|\frac{\widetilde{T}_{j,2}(k)}{\widetilde{T}_{j,1}(k)}+\frac{\widetilde{T}_{j,3}(k)}{\widetilde{T}_{j,1}(k)}+\frac{\widetilde{T}_{j,4}(k)}{\widetilde{T}_{j,1}(k)}\right|\geq 1,\,\mathcal{S}_{T}\cap\widetilde{\mathcal{S}}_{T}\right)$
$\leq$  $\,2\mathsf{P}\left(\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\frac{8\sqrt{2\gamma}\|\mathbf{N}(k)-\mathbf{N}(k_{j})\|}{d_{j}|k-k_{j}|}\geq\frac{1}{3}\right)$
$+2\mathsf{P}\left(\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\frac{4\|\mathbf{N}(k)+\mathbf{N}(k_{j})\|}{d_{j}\sqrt{2\gamma}}\geq\frac{1}{3}\right).$
By (A.59) and Assumption 7,
$\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\frac{\|\mathbf{N}(k)+\mathbf{N}(k_{j})\|}{d_{j}\sqrt{2\gamma}}\leq\max_{\gamma\leq k\leq T-\gamma}\frac{2\|\mathbf{N}(k)\|}{d_{j}\sqrt{2\gamma}}=O_{P}\left(\frac{\sqrt{\log(T/\gamma)}}{d_{j}\sqrt{2\gamma}}\right)=o_{P}(1).$
Also, we have
$\sqrt{2\gamma}\left\|\mathbf{N}(k)-\mathbf{N}(k_{j})\right\|\leq 2\left\|\sum_{t=k+1}^{k_{j}}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)\right\|\\ +\left\|\sum_{t=k-\gamma+1}^{k_{j}-\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)\right\|+\left\|\sum_{t=k+\gamma+1}^{k_{j}+\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)\right\|.$
Recalling the definition of $\mathcal{M}_{T}^{(\ell)}$ from Lemma A.13, we have
$\mathsf{P}\left(\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\frac{8\sqrt{2\gamma}\|\mathbf{N}(k)-\mathbf{N}(k_{j})\|}{d_{j}|k-k_{j}|}\geq\frac{1}{3}\right)$
$=$ $\,\mathsf{P}\left(\max_{\ell\in\{0,\pm 1\}}\max_{k_{j}-\gamma+1\leq k\leq k_{j}-Cd_{j}^{2}\omega_{T}^{2}}\frac{\sqrt{Cd_{j}^{-2}\omega_{T}}}{|k-k_{j}|}\left\|\sum_{t=k+\ell\gamma+1}^{k_{j}+\ell\gamma}\mathsf{Vech}\left(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top}-\mathbf{H}^{\top}\mathsf{E}(\mathbf{g}_{t}\mathbf{g}_{t}^{\top})\mathbf{H}\right)\right\|\geq\frac{\sqrt{C}\omega_{T}}{96}\right)$
$\leq$ $\,\mathsf{P}\left(\cap_{\ell\in\{0,\pm 1\}}\mathcal{M}_{T}^{(\ell)}\right)+o(1)=o(1),$
for large enough $C$. Altogether, we have the RHS of (A.63) bounded as $o(1)$. Analogous arguments apply to the case where $\widehat{k}_{j}>k_{j}$. Finally, setting $\omega_{T}^{(2)}=C\omega_{T}^{2}$ concludes the proof. ∎